### v1

# Supersingular curves via the Shimura–Taniyama method

Abstract.
For a curve which admits an abelian cover of the projective line branched at three points, we study when its reduction to positive characteristic is supersingular. Using the method of Shimura and Taniyama, we give a complete classification when the genus of the curve is at most 10. The natural density of the set of primes for which this construction yields a supersingular curve is larger than expected.
2020 MSC primary: 11G20, 11M38, 14G10, 14H10, 14H40. Secondary 11G10, 14G15, 14H37, 14H30.
Pries was supported by NSF grant DMS-22-00418. We thank the organizers of the RIMS conference on Theory and Applications of Supersingular Curves and Supersingular Abelian Varieties, II. We thank Everett Howe, Wanlin Li, and Elena Mantovan for helpful conversations.
1. Introduction
Suppose $k$ is an algebraically closed field of prime characteristic $p$. An elliptic curve over $k$ is supersingular if it has no points of order $p$. A smooth projective curve $X$ of genus $g$ over $k$ is supersingular if its Jacobian is isogenous to a product of $g$ supersingular elliptic curves.
One good method for constructing supersingular curves is to study a curve $X$ that is an abelian cover of the projective line $\mathbf{P}^{1}$ branched at exactly three points. The degree and ramification indices of the cover determine the genus $g$ of $X$. Such curves are quotients of a Fermat curve and their Jacobians have complex multiplication. This makes it easier to study the action of Frobenius on the cohomology [Yui80] and hence determine whether the reduction of $X$ modulo $p$ is supersingular.
Recently, we were curious which pairs $(g,p)$ occur for supersingular curves in this context. Our motivation was to show that a new construction of supersingular curves of genus $5$ [BP] applied for infinitely many new primes. We wrote this paper because we could not find a reference that states what can be determined using this method.
Let $\mathbf{S}_{g}$ be the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at exactly three points. In this paper, we determine $\mathbf{S}_{g}$ for $5\leq g\leq 10$. The main result is Theorem 6.1, which builds on data found in Sections 3, 4 and 5. We summarize it as follows.
Corollary 1.1.
For $5\leq g\leq 10$, here is the natural density $\delta_{g}$ of the set $\mathbf{S}_{g}$:
[…]
The cases $g=9,10$ are especially interesting. The codimension of the supersingular locus in the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties of dimension $g$ is $g(g+1)/2-\lfloor g^{2}/4\rfloor$. For $g\geq 9$, this exceeds the dimension of the moduli space $\mathcal{M}_{g}$ of smooth curves of genus $g$. For this reason, the existence of a smooth supersingular curve of genus $g$ is geometrically unexpected for $g\geq 9$.
Work of Yui [Yui80] implies that $\delta_{g}$ is positive for all $g$, which can also be deduced from our analysis, see Remark 5.3.
For the proof, we use Ekedahl’s work on [Eke87, page 173] to classify the degrees and ramification indices of abelian covers $X\to\mathbf{P}^{1}$ branched at 3 points with fixed genus $g$. When $g=4,5$, Ekedahl determines conditions on $p$ for which $X$ is superspecial (meaning its Jacobian is isomorphic to a product of supersingular elliptic curves). We work instead with the Shimura–Taniyama formula [Tat71, Section 5], a method used also in [LMPT19b], to determine the set of primes $p$ for which the reduction is supersingular. We use SageMath to enumerate all the possibilities for genus $5\leq g\leq 10$.
We restrict to the case $g\geq 5$ and $p$ odd because of earlier work: for $g=1$ by Deuring [Deu41]; for $g=2$ by Serre [Ser83, Théorème 3], also [IKO86, Proposition 3.1]; for $g=3$ by Oort [Oor91, Theorem 5.12], also [Ibu93, Theorem 1]; and for $g=4$ by Harashita, Kudo, and Senda [KHS20, Corollary 1.2], also [Pri24, Theorem 1.1]. For $p=2$, for all $g$, see the work of van der Geer and van der Vlugt [vdGvdV95].
2. Abelian covers branched at three points
Let $X$ be a smooth projective connected curve over $k$. Let $\phi:X\to{\mathbf{P}}^{1}$ be a Galois cover branched exactly at $B:=\{0,1,\infty\}$. Let $g$ be the genus of $X$. We suppose that the Galois group $G$ of $\phi$ is abelian and that $\phi$ is tamely ramified.
2.1. The degree and ramification indices
For $b\in B$, let $I_{b}$ be the inertia group above $b$ and let $c_{b}=|I_{b}|$. Since $\phi$ is tamely ramified, $I_{b}$ is cyclic. Without loss of generality, we suppose that $c_{0}\geq c_{1}\geq c_{\infty}$.
Definition 2.1.
The ramification type for the cover $\phi$ is the tuple $z=[c_{0},c_{1},c_{\infty},s]$, where $s$ is the index of $I_{0}$ in $G$.
The ramification type $z$ is essentially what Ekedahl studied, except that we switch the role of 00 and $\infty$ to simplify the equation for $\phi$.
Notation 2.2.
Let $r=\mathrm{gcd}(c_{0},c_{1},c_{\infty})$. For distinct $i,j\in B$, let $e_{i,j}=\mathrm{gcd}(c_{i},c_{j})/r$.
Any two of $e_{0,1}$, $e_{0,\infty}$, and $e_{1,\infty}$ are relatively prime and $e_{0,1}\geq e_{0,\infty}\geq e_{1,\infty}$.
Lemma 2.3.
[Eke87, page 173] With notation as above:
(2.1) $2g-2=s(re_{0,1}e_{0,\infty}-(1+(e_{0,1}/e_{1,\infty})+(e_{0,\infty}/e_{1,\infty})).$
Proof.
Above $b\in B$, the number of ramification points is #$\#G/c_{b}$ and each has inertia group of order $c_{b}$. By definition, $|G|=sc_{0}$. By the Riemann–Hurwitz formula:
$2\displaystyle 2g-$ 2 $=$ $-2|G|+\sum_{b\in B}(|G|/c_{b})(c_{b}-1)$
$=$ $|G|(1-((1/c_{0})+(1/c_{1})+(1/c_{\infty})))$
$=$ $s(c_{0}-(1+(c_{0}/c_{1})+(c_{0}/c_{\infty}))).$
The result follows since $c_{0}=re_{0,1}e_{0,\infty}$, $c_{1}=re_{0,1}e_{1,\infty}$, and $c_{\infty}=re_{0,\infty}e_{1,\infty}$. ∎
We identify $G$ with $\mathrm{Gal}(\phi)$ and define a refinement of the ramification type.
Definition 2.4.
See e.g., [Bou01, Definitions 3.2 & 4.1]. The inertia type $a$ is the triple $a=(a_{0},a_{1},a_{\infty})$, where $a_{b}$ denotes the canonical generator for $I_{b}$, found from the action of the inertia group on a local parameter at a ramification point above $b$.
There are three constraints on the inertia type: $|a_{b}|=c_{b}$ for $b\in B$; $a_{0},a_{1},a_{\infty}$ generate $G$; and $a_{0}a_{1}a_{\infty}=\mathrm{Id}_{G}$. This implies that $s$ divides $c_{1}$ and $s$ divides $c_{\infty}$. Not every ramification type $z=[c_{0},c_{1},c_{\infty},s]$ has an inertia type that satisfies these constraints. See Remark 5.1 for some examples.
Given $z$, we can change the labeling of branch points whose inertia groups have the same order; this allows us to permute these entries of the inertia type. Also, we can change the identification of $G$ with $\mathrm{Gal}(\phi)$; this modifies each entry of the inertia type by the action of some $\sigma\in\mathrm{Aut}(G)$. Two inertia types are equivalent, denoted $\approx$, if they are in the same orbit under these actions.
2.2. Cyclic covers branched at three points
In this section, we suppose $\phi$ is cyclic of order $m$, and identify $G=\mathbf{Z}/m\mathbf{Z}$. The inertia type $(a_{0},a_{1},a_{\infty})$ is defined up to multiplication by $(\mathbf{Z}/m\mathbf{Z})^{*}$. Note that $c_{b}=|a_{b}|=m/\gcd(m,a_{b})$. An affine equation for the cover is
$X:y^{m}=x^{a_{0}}(1-x)^{a_{1}},$
with $G$ acting by multiplication by $m$th roots of unity on $y$.
Definition 2.5.
For $1\leq j<m$, let $f_{j}$ be the dimension of the $\zeta_{m}^{j}$-eigenspace of $H^{0}(X,\Omega^{1}_{X})$. The signature type of $\phi$ is $f=(f_{1},\ldots,f_{m-1})$.
Let $\langle q\rangle$ denote the fractional part of $q\in\mathbf{Q}$. Then $f_{j}=-1+\sum_{b\in B}\langle\frac{-ja_{b}}{m}\rangle$ by e.g., [Moo10, Lemma 2.7, Section 3.2]. For degree $m$ cyclic covers $\phi:X\to{\mathbf{P}}^{1}$ branched at $B=\{0,1,\infty\}$, the following facts are well-known (e.g., [LMPT19b, Lemma 3.1]):
  • the Jacobian ${\mathbf{Jac}}(X)$ has complex multiplication by $\prod_{d}\mathbf{Q}(\zeta_{d})$, where the product is taken over all integers $d$ such that $1<d\mid m$ and $d\nmid a_{b}$ for any $b\in B$;
  • if $1\leq j<m$, then $f_{j}\in\{0,1\}$, and $f(j)+f(m-j)=1$ provided $d_{j}$, the order of $j$ modulo $m$, does not divide $a_{b}$ for any $b\in B$.
2.3. Non-cyclic Galois groups and inertia types
The following determines the Galois group in all the cases with $s>1$ for $5\leq g\leq 10$.
Lemma 2.6.
Let $z=[c_{0},c_{1},c_{\infty},s]$ be a ramification type such that $s>1$, $s\mid c_{\infty}$, $c_{\infty}\mid c_{1}$, and $c_{1}\mid c_{0}$. Then $G$ is non-cyclic with $G\simeq\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$.
Proof.
By hypothesis, there is a prime $\ell$ dividing all three of $s=c_{0}s/c_{0}$, $c_{0}s/c_{1}$, and $c_{0}s/c_{\infty}$. If $G$ was cyclic, $\ell$ would divide each of $m,a_{0},a_{1},a_{\infty}$, contradicting the fact that $a_{0},a_{1},a_{\infty}$ generate $G$. Since $G$ is generated by two elements $a_{0},a_{1}$, this implies that $G\simeq\mathbf{Z}/r_{1}\mathbf{Z}\times\mathbf{Z}/r_{2}\mathbf{Z}$ where $r_{1}r_{2}=c_{0}s$ and $r_{2}\mid r_{1}$. Note that $r_{1}$ is a multiple of $c_{0}$ since $a_{0}$ has order $c_{0}$. The hypothesis on $z$ implies that the orders of $a_{1}$ and $a_{\infty}$ divide $c_{0}$. So $a_{0}$ and $a_{1}$ generate $G$ only if $r_{1}=c_{0}$. Thus $G\simeq\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$. ∎
For $s>1$, we denote elements of $G=\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$ as an ordered pair. So each $a_{b}$ in the inertia type $a=(a_{0},a_{1},a_{\infty})$ is written as $a_{b}=(a_{b,1},a_{b,2})$. To choose a representative for the equivalence class of $a$, we take the natural inclusion of $G$ in $V=(\mathbf{Z}/c_{0}\mathbf{Z})^{2}$, and use that $\mathrm{Aut}(V)$ acts transitively on pairs of linearly independent elements of order $c_{0}$. Without loss of generality, we choose $a_{0}=(1,0)$. When $c_{\infty}=s$, we also choose $a_{\infty}=(0,1)$.
2.4. Newton polygons for cyclic covers branched at three points
Suppose $p$ is a prime not dividing $m$, the degree of the cyclic cover $\phi$. We consider the reductions of $X$ and ${\mathbf{Jac}}(X)$ modulo $p$. The Newton polygon of ${\mathbf{Jac}}(X)$ is an invariant that measures the $p$-adic valuations of the eigenvalues of Frobenius. In this section, we explain how to compute the Newton polygon from the signature type of $\phi$.
Definition 2.7.
For non-negative integers $c$ and $d$ with ${\mathrm{gcd}}(c,d)=1$, fix a $p$-divisible group $G_{c,d}$ of codimension $c$, dimension $d$, and thus height $c+d$. The slope of $G_{c,d}$ is $\lambda=d/(c+d)$, with multiplicity $c+d$. Let $G_{\lambda}=G_{c,d}$.
Suppose $A$ is a principally polarized abelian variety of dimension $g$ over $k$, By the Dieudonné–Manin classification, the $p$-divisible group $A[p^{\infty}]$ is isogenous to $\oplus_{\lambda}G_{\lambda}^{m_{\lambda}}$ for some multi-set of non-negative integers $m_{\lambda}$. This multi-set defines a symmetric Newton polygon $\nu(A)$ of height $2g$ with integer break-points.
Remark 2.8.
Suppose $A$ is defined over $\mathbf{F}_{0}$, a finite field of characteristic $p$. Let $W(\mathbf{F}_{0})$ denote the Witt vector ring of $\mathbf{F}_{0}$. Consider the action of Frobenius $\varphi$ on the crystalline cohomology group $H^{1}_{\rm cris}(A/W(\mathbf{F}_{0}))$. There exists an integer $n$ such that $\varphi^{n}$, the composite of $n$ Frobenius actions, is a linear map. The Newton polygon $\nu(A)$ is the multi-set of rational numbers $\lambda$ such that $n\lambda$ are the valuations at $p$ of the eigenvalues of Frobenius for this action.
We sometimes describe the Newton polygon of a curve $X$ in terms of the decomposition of ${\mathbf{Jac}}(X)[p^{\infty}]$. For example, a curve of genus $g$ is ordinary if its $p$-divisible group is $(G_{0,1}\oplus G_{1,0})^{g}$ and is supersingular if its $p$-divisible group is $G_{1,1}^{g}$.
Notation 2.9.
Let $R$ denote the set of orbits of the multiplication by $p$ map [×$[\times p]$ on the set $\{j\in\mathbf{Z}/m\mathbf{Z}:j\neq 0\}$. For each orbit $r\in R$, the additive order $d_{j}$ of $j$ modulo $m$ is the same for all values $j$ in $r$ and we denote it by $d_{r}$. For $\epsilon\in\{0,1\}$, let
[…]
For $r\in R$ such that $d_{r}$ does not divide $a_{b}$ for $b\in B$, set $\alpha_{r}=\#(r\cap S_{1})$.
The set $R$ is in one-to-one correspondence with the set of primes of the group ring $\mathbf{Q}[\mu_{m}]$ above $p$. The $p$-divisible group ${\mathbf{Jac}}(X)[p^{\infty}]$ canonically decomposes into pieces ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ indexed by the orbits $r\in R$.
By the Shimura–Taniyama formula [Tat71, Section 5], the only slope of the Newton polygon of ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ is $\lambda_{r}:=\alpha_{r}/\#r$. In particular, if every orbit is self-dual, meaning invariant under negation modulo $m$, then $X$ is supersingular.
Example 2.10.
Let $m=8$ and $p\equiv 3\bmod 4$. The orbits are $r_{1}=\{1,3\}$, $r_{2}=\{2,6\}$, $r_{5}=\{5,7\}$, and $r_{4}=\{4\}$.
If $a=(1,1,6)$, then $\phi$ has signature $f=(1,1,1,0,0,0,0)$. Thus $S_{1}=\{1,2,3\}$. Then $\lambda_{r_{1}}=1$, $\lambda_{r_{2}}=1/2$, and $\lambda{r_{5}}=0$. Thus $\nu(\mathbf{Jac}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$.
If $a=(1,5,2)$, then $\phi$ has signature $f=(1,1,0,0,1,0,0)$. Thus $S_{1}=\{1,2,5\}$ and $X$ is supersingular.
Remark 2.11.
One can also determine whether $X$ is supersingular using a method of Yui [Yui80] and Aoki [Aok08] that relies on $p$-adic valuations of Jacobi sums. By [Aok08, Theorem 1.1], $X$ is supersingular if and only if either (i) $p^{i}\equiv-1\bmod m$ for some $i$; or (ii) $a\approx(1,-p^{i},p^{i}-1)$ for some $i$ such that $d=\mathrm{gcd}(p^{i}-1,m)>1$ and $p^{j}\equiv-1\bmod m/d$ for some $j$.
2.5. Decomposition from the Kani–Rosen theorem
Suppose $\phi:X\to\mathbf{P}^{1}$ is an abelian non-cyclic cover branched at $B=\{0,1,\infty\}$. It is possible to generalize Sections 2.2 and 2.4 to the abelian non-cyclic case. This would take us too far afield so we use the Kani–Rosen theorem.
The goal is to decompose $\mathrm{Jac}(X)$, up to isogeny, as the product $\prod_{i\in I}J_{i}$ of abelian varieties $J_{i}$ of smaller dimension. Each abelian variety $J_{i}$ in this decomposition is the Jacobian of a curve $Z_{i}$, such that $Z_{i}$ is a quotient of $X$ by a subgroup $H_{i}\subset G$.
Theorem 2.12 (Kani–Rosen [KR89, Theorem B]).
Let $G,H_{1},\ldots,H_{t}\subset\mathrm{Aut}(X)$ be such that $G=\cup_{i=1}^{t}H_{i}$ and $H_{i}\cap H_{j}=\{\mathrm{id}\}$ for $i\not=j$. Let $n_{G}=\#G$ and $n_{i}=\#H_{i}$ for $1\leq i\leq t$. Then there is an isogeny:
(2.3) $\mathrm{Jac}(X)^{t-1}\times\mathrm{Jac}(X/G)^{n_{G}}\sim\prod_{i=1}^{t}\mathrm{Jac}(X/H_{i})^{n_{i}}.$
Suppose $G\simeq(\mathbf{Z}/2\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3}\subset G$ be the three subgroups of order two. Then (2.3) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{2}\sim\prod_{i=1}^{3}\mathrm{Jac}(X/H_{i})$. Suppose $G\simeq(\mathbf{Z}/3\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3},H_{4}\subset G$ be the four subgroups of order three. Then (2.3) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{3}\sim\prod_{i=1}^{4}\mathrm{Jac}(X/H_{i})$.
2.6. Enumerating and analyzing abelian covers
Using SageMath, we automated the search for ramification data producing abelian covers $\phi:X\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ whose genus lies in a specified range. We outline the approach, which depends on whether $s=1$ and $s>1$.
When $s=1$, then $\phi$ is cyclic and the program determines possible inertia types. Without loss of generality, we identified the generator of $I_{0}$ with $1\in G$ and so $a_{0}=1$. Then the program analyzes the congruence classes of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular, using the Shimura–Taniyama formula as in Section 2.4.
For each case when $s>1$, Lemma 2.6 implies that $G$ is non-cyclic and that $G\simeq(\mathbf{Z}/c_{0}\mathbf{Z})\times(\mathbf{Z}/s\mathbf{Z})$. For every $z$, we determined the possible inertia types $a$. Without loss of generality, we identified the generator of $I_{0}$ with $a_{0}=(1,0)$. In the small number of non-cyclic cases with $5\leq g\leq 10$, we determined by hand a decomposition of $\mathbf{Jac}(X)$ into Jacobians of curves that are cyclic covers of $\mathbf{P}^{1}$ branched at $B$, using the Kani–Rosen method as in Section 2.5. This allowed us to find the supersingular primes.
In Section 3, we give a complete analysis of the genus $5$ situation to illustrate the methods. In particular, abelian covers of $\mathbf{P}^{1}$ branched at three points do not produce supersingular curves of genus $5$ for every prime $p\equiv 3\bmod 4$. Section 4 contains the results of our program [BP25] for cyclic cases when $6\leq g\leq 10$. Section 5 treats the non-cyclic cases when $6\leq g\leq 10$.
3. Refining Ekedahl’s Results for genus five
Suppose $\phi:X\to\mathbf{P}^{1}$ is a tamely ramified abelian cover branched at $B=\{0,1,\infty\}$. In [Eke87, page 173], Ekedahl enumerates ways $\phi$ can ramify so that $X$ has genus $5$. There are seven non-equivalent cases, as confirmed by our program, which are listed in Propositions 3.1, 3.3, and 3.6.
Ekedahl studies the situation when $X$ is superspecial, meaning that $\mathrm{Jac}(X)$ is isomorphic to a product of supersingular elliptic curves. The superspecial condition is equivalent to the Frobenius operator being trivial on $H^{1}(X,\mathcal{O}_{X})$. For each of the seven cases, Ekedahl determines the primes $p$ for which $X$ is superspecial, thus proving that there exists a superspecial smooth curve of genus $5$ which is an abelian cover of $\mathbf{P}^{1}$ branched at $3$ points if and only if $p\equiv-1\bmod 11,12,15,20,8$ or $p\equiv 11\bmod 15$.
If $X$ is superspecial, then it is supersingular, but the converse is not always true. To determine the (less restrictive) conditions on $p$ for which $X$ is supersingular, we use the Shimura–Taniyama formula.
3.1. Cyclic cases for genus five
We use the abbreviation QNR for quadratic non-residue.
Proposition 3.1.
If $\phi:X\to\mathbf{P}^{1}$ is a cyclic cover branched at $3$ points and $X$ has genus $5$, then the ramification type $z$ for $\phi$ is one of the following:
  1. (1)
    $z=[11,11,11,1]$. Then $X$ is supersingular if and only if $p$ is a QNR modulo $11$.
  2. (2)
    $z=[15,15,3,1]$. Then $X$ is supersingular if and only if $p\equiv-1,11\bmod 15$.
  3. (3)
    $z=[22,11,2,1]$. Then $X$ is supersingular if and only if $p$ is a QNR modulo $11$.
  4. (4)
    $z=[12,12,6,1]$. Then $X$ is supersingular if and only if $p\equiv-1\bmod 12$.
  5. (5)
    $z=[20,20,2,1]$. Then $X$ is supersingular if and only if $p\equiv-1,11\bmod 20$.
Proof.
By [Eke87, page 173], these are the only options for $z$ when $\phi$ is cyclic and $X$ has genus $5$. We analyze these covers the same way our program does to illustrate how it works. In each case, $G=\mathbf{Z}/m\mathbf{Z}$ with $m=c_{0}$ and $\phi$ is totally ramified over $b=0$. A short calculation determines the choices for the inertia type $a$, with $a_{0}=1$.
Recall that $R$ denotes the set of orbits of ×$\times p$ on $\mathbf{Z}/m\mathbf{Z}$. In each situation, if $p^{i}\equiv-1\bmod m$ for some $i$, then each orbit is self-dual and so $X$ is supersingular. When $m$ is composite, there can be additional cases where $\alpha_{r}/\#r=1/2$ for each $r\in R$. We give some additional details in each case.
  1. (1)
    Note that $a\approx(1,k,-(1+k))$ for some $1\leq k\leq 9$. The result is proven as one case of [LMPT19b, Theorem 5.1].
  2. (2)
    Note that $a\approx(1,4,10)$. Then $S_{1}=\{1,2,4,5,8\}$. When $p\equiv 11\bmod 15$, then $X$ is supersingular. This can be verified using the Shimura–Taniyama formula because #$\#r=2$ for each $r\in R$ and each orbit contains exactly one element of $S_{1}$. The only other case when $X$ is supersingular is when $p\equiv-1\bmod{15}$.
  3. (3)
    Then $a\approx(1,10,11)$. The condition $p^{i}\equiv-1\bmod 22$ is solvable for some $i$ if and only if $p$ is a quadratic non-residue modulo $11$.
  4. (4)
    Then $a\approx(1,1,10)$ and $S_{1}=\{1,2,3,4,5\}$. The only case when $X$ is supersingular is when $p\equiv-1\bmod 12$.
  5. (5)
    Then $a\approx(1,9,10)$ and $S_{1}=\{1,3,5,7,9\}$. When $p\equiv 11\bmod 20$, then #$\#r=2$ for each $r\in R$ and each orbit contains exactly one element of $S_{1}$. By the Shimura–Taniyama formula, $X$ is supersingular (but not superspecial). ∎
3.2. Non-cyclic cases for genus five
Suppose $\phi:X\to\mathbf{P}^{1}$ is an abelian non-cyclic cover branched at $B=\{0,1,\infty\}$ and $X$ has genus $5$. Ekedahl shows there are two options for the ramification type: either (6) $z=[8,8,4,2]$ or (7) $z=[12,12,2,2]$. In both cases, we determine the congruence conditions on $p$ for which $X$ is supersingular.
3.2.1. Case (6)
When $z=[8,8,4,2]$, then $G$ has order $16$. By Lemma 2.6, $G\simeq\mathbf{Z}/8\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. Let $H\subset G$ be the subgroup generated by $(4,0)$ and $(0,1)$. Let $E_{1728}$ be the elliptic curve with $j$-invariant $1728$, with affine equation $y^{2}=x^{3}-x$.
Lemma 3.2.
When $z=[8,8,4,2]$, then:
  1. (1)
    the quotient of $X$ by $H$ is isomorphic to $E_{1728}$;
  2. (2)
    the $H$-Galois subcover $X\to E_{1728}$ factors through three curves $E^{\prime}$, $W$ and $W^{\prime}$, where $E^{\prime}\to E_{1728}$ is a $2$-isogeny, and $W$ and $W^{\prime}$ have genus $3$; and
  3. (3)
    there is an isogeny ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})\sim{\mathbf{Jac}}(W)\times{\mathbf{Jac}}(W^{\prime})$.
Proof.
When $z=[8,8,4,2]$, then $G\simeq\mathbf{Z}/8\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. After modifying by an element of $\mathbf{Aut}(G)$, the inertia type can be chosen to be $a=((1,0),(1,1),(6,1))$.
The $G$-cover $X\to{\mathbf{P}}^{1}$ has: $2$ points above 00 with $I_{0}=\langle(1,0)\rangle$; $2$ points above $1$ with $I_{1}=\langle(1,1)\rangle$; and $4$ points above $\infty$ with $I_{\infty}=\langle(6,1)\rangle$.
  1. (1)
    Note that $G/H\simeq\mathbf{Z}/4\mathbf{Z}$. So $\psi:X/H\to{\mathbf{P}}^{1}$ is a $\mathbf{Z}/4\mathbf{Z}$-cover branched at $B$. The intersection of $H$ with each of $I_{0}$, $I_{1}$, and $I_{\infty}$ has order $2$. By the Riemann–Hurwitz formula, $X/H$ has genus $1$. Since $\psi$ is totally ramified above 00 and $\infty$, this implies that $X/H$ is isomorphic to $E_{1728}$.
  2. (2)
    Let $H_{1}:=\langle(4,0)\rangle$. Since $(4,0)$ is in the inertia group at every point of $X$ above 00, $1$, and $\infty$, the double cover $X\to X/{H_{1}}$ is branched at $8$ points. By the Riemann–Hurwitz formula, $X/{H_{1}}$ is an elliptic curve, which we denote $E^{\prime}$. The double cover $E^{\prime}\to E_{1728}$ is unramified and is thus a $2$-isogeny.
    Let $W$ be the quotient of $X$ by $H_{2}:=\langle(0,1)\rangle$. Since $H_{2}\cap I_{b}$ is trivial for each $b\in B$, the double cover $X\to W$ is unramified. By the Riemann–Hurwitz formula, $W$ has genus $3$. The same holds for the quotient $W^{\prime}$ of $X$ by $H_{3}:=\langle(4,1)\rangle$.
  3. (3)
    By Theorem 2.12, ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})^{2}\sim{\mathbf{Jac}}(E^{\prime})\times{\mathbf{Jac}}(W)\times{\mathbf{Jac}}(W^{\prime})$. The result follows since the category of abelian varieties up to isogeny is semi-simple, and ${\mathbf{Jac}}(E_{1728})\sim{\mathbf{Jac}}(E^{\prime})$. ∎
Proposition 3.3.
Suppose $z=[8,8,4,2]$. Then $X$ is supersingular if and only if $p\equiv 7\bmod 8$. If $p\equiv 3\bmod 8$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
Proof.
By Lemma 3.2(i), $E_{1728}$ is a quotient of $X$. We restrict to the case $p\equiv 3\bmod 4$ which is equivalent to $E_{1728}\sim E^{\prime}$ being supersingular. In this case, $X$ is supersingular if and only if $W$ and $W^{\prime}$ are both supersingular, by Lemma 3.2(iii).
The cover $W\to\mathbf{P}^{1}$ has Galois group $\mathbf{Z}/8\mathbf{Z}$ and is branched above $B$ with inertia type $\overline{a}=(1,1,6)$. The cover $W^{\prime}\to\mathbf{P}^{1}$ has Galois group $\overline{G}=G/\langle(4,1)\rangle\simeq\mathbf{Z}/8\mathbf{Z}$, which is generated by $z=(1,1)$. We compute its inertia type $\overline{a}^{\prime}=(5z,z,2z)\approx(1,5,2)$.
If $p\equiv 7\bmod 8$, then every orbit of ×$\times p$ on $\{1\leq j<8\}$ is self-dual, and so both $W$ and $W^{\prime}$ are supersingular by the Shimura–Taniyama formula. Suppose $p\equiv 3\bmod 8$. By Example 2.10, the curve $W^{\prime}$ is supersingular since $\overline{a}^{\prime}\approx(1,5,2)$, while the curve $W$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$ since $\overline{a}=(1,1,6)$. Thus ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})$ has two slopes of 00 and $1$, and $X$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$. ∎
Remark 3.4.
Another construction of supersingular curves of genus $5$ when $p\equiv 7\bmod 8$ and $p\gg 0$ can be found in [LMPT19a, Theorem 1.2].
3.2.2. Case (7)
In case (7), $z=[12,12,2,2]$. Recall that $E_{1728}$ is the elliptic curve with $j$-invariant $1728$. Let $E_{0}$ be the elliptic curve with $j$-invariant 00, with affine equation $y^{2}=x^{3}-1$.
Lemma 3.5.
When $z=[12,12,2,2]$, then:
  1. (1)
    the curve $X$ is a Klein-four cover of $\mathbf{P}^{1}$, whose degree $2$ quotients $Z$, $W$, and $P$ have genera $3$, $2$, and 00 respectively, and there is an isogeny ${\mathbf{Jac}}(X)\sim{\mathbf{Jac}}(Z)\times{\mathbf{Jac}}(W)$;
  2. (2)
    the curve $Z$ is a $\mathbf{Z}/12\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at three points, with inertia type $(1,5,6)$ and $Z$ is a $\mathbf{Z}/3\mathbf{Z}$-cover of the elliptic curve $E_{1728}$;
  3. (3)
    the curve $W$ is a Klein-four cover of $\mathbf{P}^{1}$, whose three degree $2$ quotients are isomorphic to $E_{0}$, $E_{0}$, and $\mathbf{P}^{1}$, and there is an isogeny ${\mathbf{Jac}}(W)\sim E_{0}^{2}$.
Proof.
By Lemma 2.6, $G\simeq\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. After modifying by an element of $\mathbf{Aut}(G)$, we can suppose that $a=((1,0),(11,1),(0,1))$.
  1. (1)
    The automorphism $(0,1)$ fixes the twelve points of $X$ above $\infty$. So the quotient $P$ of $X$ by $I_{\infty}$ is $\mathbf{P}^{1}$. The automorphism $(6,0)$ fixes the two points of $X$ above 00 and the two points of $X$ above $1$. So the quotient $W$ of $X$ by $\langle(6,0)\rangle$ has genus $2$. The automorphism $(6,1)$ fixes no points of $X$. So the quotient $Z$ of $X$ by $\langle(6,1)\rangle$ has genus $3$. By Theorem 2.12, ${\mathbf{Jac}}(X)\sim{\mathbf{Jac}}(Z)\times{\mathbf{Jac}}(W)$.
  2. (2)
    The curve $Z$ has an action by $G/\langle(6,1)\rangle\simeq\mathbf{Z}/12\mathbf{Z}$, which is generated by $w=(1,1)$. The inertia type $a=((1,0),(11,1),(0,1))$ reduces to $\overline{a}=(7w,11w,6w)\approx(5,1,6)$. There is an order $3$ automorphism $\sigma$ on $Z$ that fixes two points. The quotient of $Z$ by $\sigma$ is an elliptic curve with an automorphism of order $4$ that fixes two points. Thus $Z$ is a $\mathbf{Z}/3\mathbf{Z}$-cover of $E_{1728}$.
  3. (3)
    The curve $W$ admits an action of $\mathbf{Z}/6\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. Each of the automorphisms $(3,0)$ and $(3,1)$ fixes exactly two points of $W$, so $W$ admits two double covers of an elliptic curve. In both cases, this elliptic curve is $E_{0}$ because it has an automorphism of order $3$ which fixes three points. Since $(3,0)$ and $(3,1)$ generate a Klein-four group, ${\mathbf{Jac}}(W)\sim E_{0}^{2}$ by Theorem 2.12. ∎
Proposition 3.6.
Suppose $z=[12,12,2,2]$.
Then $X$ is supersingular if and only if $p\equiv 11\bmod 12$.
If $p\equiv 5\bmod 12$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{3}\oplus G_{1,1}^{2}$.
If $p\equiv 7\bmod 12$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
Proof.
Note that $E_{0}$ is supersingular exactly when $p\equiv 2\bmod 3$ and $E_{1728}$ is supersingular exactly when $p\equiv 3\bmod 4$. By Lemma 3.5, $X$ dominates $E_{0}$ and $E_{1728}$. Thus $X$ supersingular implies that $p\equiv 11\bmod 12$.
Suppose $p\equiv 11\bmod 12$. Then $E_{0}$ is supersingular. By Lemma 3.5, ${\mathbf{Jac}}(X)\sim{\mathbf{Jac}}(Z)\times E_{0}^{2}$. The curve $Z$ is supersingular by the Shimura–Taniyama formula since the orbits of ×$\times p$ on $\mathbf{Z}/12\mathbf{Z}$ are self-dual. This completes the proof of the first claim.
The $\mathbf{Z}/12\mathbf{Z}$-cover $Z\to\mathbf{P}^{1}$ has inertia type $(5,1,6)$. We compute that the eigenspaces of $H^{0}(Z,\Omega^{1}_{Z})$ of positive dimension are indexed by $S_{1}=\{1,3,5\}$. When $p\equiv 5\bmod 12$, then $E_{0}$ is supersingular. Also the orbits of ×$\times p$ are $\{1,5\}$, $\{7,11\}$, $\{3\}$ and $\{9\}$. By the Shimura–Taniyama formula, the Newton polygon of $Z$ is $(G_{0,1}\oplus G_{1,0})^{3}$. When $p\equiv 7\bmod 12$, then $E_{0}$ is ordinary. Also the orbits of ×$\times p$ are $\{1,7\}$, $\{5,11\}$, and $\{3,9\}$. By the Shimura–Taniyama formula, $Z$ is supersingular. ∎
4. Cyclic Examples Genus 6 through 10
Our program computed the following data for curves $X$, of genus $6\leq g\leq 10$, that admit a degree $m$ cyclic cover $\phi:X\to\mathbf{P}^{1}$ branched at three points: the ramification type $z$; the possible inertia types $a$, up to equivalence; and the odd primes $p$ with $\mathrm{gcd}(p,m)=1$ such that $X$ is supersingular. The results are listed in Tables 1-5
$z$ $a$ supersingular primes
$\left[13,13,13,1\right]$ $\left(1,11,1\right),\left(1,10,2\right),\left(1,9,3\right)$ $p\not\equiv 1,3,9\bmod{13}$
$\left[14,14,7,1\right]$ $\left(1,11,2\right),\left(1,5,8\right),\left(1,1,12\right)$ $p\equiv QNR\bmod{14}$
$\left[15,15,5,1\right]$ $\left(1,11,3\right)$ $p\equiv 2,3,4\bmod 5$
$\left(1,8,6\right)$ $p\equiv 7,13,14\bmod{15}$
$\left[16,16,4,1\right]$ $\left(1,11,4\right)$ $p\equiv 15\bmod{16}$
$\left[20,5,4,1\right]$ $\left(1,4,15\right)$ $p\equiv 19\bmod{20}$
$\left[18,18,3,1\right]$ $\left(1,11,6\right)$ $p\equiv 2\bmod{3}$
$\left[21,7,3,1\right]$ $\left(1,6,14\right)$ $p\equiv 5,17,20\bmod{21}$
$\left[24,24,2,1\right]$ $\left(1,11,12\right)$ $p\equiv 13,23\bmod{24}$
$\left[26,13,2,1\right]$ $\left(1,12,13\right)$ $p\not\equiv 1,3,9\bmod{26}$
Table 1. Cyclic genus $6$
$z$ $a$ supersingular primes
$\left[15,15,15,1\right]$ $\left(1,13,1\right)$ $p\equiv 14\bmod{15}$
$\left[16,16,8,1\right]$ $\left(1,13,2\right),\left(1,1,14\right)$ $p\equiv 15\bmod{16}$
$\left(1,9,6\right)$ $p\equiv 7,15\bmod{16}$
$\left[18,9,6,1\right]$ $\left(1,14,3\right),\left(1,2,15\right)$ $p\equiv 2\bmod{3}$
$\left[20,10,4,1\right]$ $\left(1,14,5\right)$ $p\equiv 19\bmod{20}$
$\left[21,21,3,1\right]$ $\left(1,13,7\right)$ $p\equiv 2\bmod{3}$
$\left[24,8,3,1\right]$ $\left(1,15,8\right)$ $p\equiv 23\bmod{24}$
$\left[28,28,2,1\right]$ $\left(1,13,14\right)$ $p\equiv 3\bmod{4}$
$\left[30,15,2,1\right]$ $\left(1,14,15\right)$ $p\equiv 29\bmod{30}$
Table 2. Cyclic genus $7$
$z$ $a$ supersingular primes
$\left[17,17,17,1\right]$ $\left(1,15,1\right),\left(1,14,2\right),\left(1,13,3\right)$ $p\not\equiv 1\bmod{17}$
$\left[18,18,9,1\right]$ $\left(1,13,4\right),\left(1,1,16\right)$ $p\equiv 2\bmod{3}$
$\left[20,20,5,1\right]$ $\left(1,11,8\right)$ $p\equiv 2,3,4\bmod{5}$
$\left(1,7,12\right)$ $p\equiv 13,17,19\bmod{20}$
$\left[24,24,3,1\right]$ $\left(1,7,16\right)$ $p\equiv 17,23\bmod{24}$
$\left[32,32,2,1\right]$ $\left(1,15,16\right)$ $p\not\equiv 1,15\bmod{32}$
$\left[34,17,2,1\right]$ $\left(1,16,17\right)$ $p\not\equiv 1\bmod{34}$
Table 3. Cyclic genus $8$
$z$ $a$ supersingular primes
$\left[19,19,19,1\right]$ $\left(1,17,1\right),\left(1,16,2\right)$, $p\equiv QNR\bmod{19}$
$\left(1,15,3\right),\left(1,11,7\right)$ $p\equiv QNR\bmod{19}$
$\left[20,20,10,1\right]$ $\left(1,17,2\right)$ $p\equiv 3,7,19\bmod{20}$
$\left(1,1,18\right)$ $p\equiv 19\bmod{20}$
$\left[21,21,7,1\right]$ $\left(1,17,3\right),\left(1,11,9\right)$ $p\equiv 5,10,17,19,20\bmod{21}$
$\left(1,8,12\right)$ $p\equiv QNR\bmod{7}$
$\left[24,8,6,1\right]$ $\left(1,3,20\right)$ $p\equiv 23\bmod{24}$
$\left[24,24,4,1\right]$ $\left(1,17,6\right)$ $p\equiv 7,23\bmod{24}$
$\left(1,5,18\right)$ $p\equiv 19,23\bmod{24}$
$\left[28,7,4,1\right]$ $\left(1,20,7\right)$ $p\equiv 3,19,27\bmod{28}$
$\left[27,27,3,1\right]$ $\left(1,17,9\right)$ $p\equiv 2\bmod{3}$
$\left[30,10,3,1\right]$ $\left(1,9,20\right)$ $p\equiv 29\bmod{30}$
$\left[36,36,2,1\right]$ $\left(1,17,18\right)$ $p\equiv 3\bmod{4}$
$\left[38,19,2,1\right]$ $\left(1,18,19\right)$ $p\equiv QNR\bmod{38}$
Table 4. Cyclic genus $9$
$z$ $a$ supersingular primes
$\left[21,21,21,1\right]$ $\left(1,19,1\right),\left(1,16,4\right)$ $p\equiv 5,17,20\bmod{21}$
$\left[22,22,11,1\right]$ $\left(1,19,2\right),\left(1,17,4\right),\left(1,15,6\right)$ $p\equiv QNR\bmod{22}$
$\left(1,9,12\right),\left(1,1,20\right)$ $p\equiv QNR\bmod{22}$
$\left[24,12,8,1\right]$ $\left(1,14,9\right),\left(1,2,21\right)$ $p\equiv 23\bmod{24}$
$\left[24,24,6,1\right]$ $\left(1,19,4\right)$ $p\equiv 5,23\bmod{24}$
$\left[25,25,5,1\right]$ $\left(1,19,5\right),\left(1,14,10\right)$ $p\neq 1\bmod{5}$
$\left[30,6,5,1\right]$ $\left(1,5,24\right)$ $p\equiv 29\bmod{30}$
$\left[28,14,4,1\right]$ $\left(1,6,21\right)$ $p\equiv 3,19,27\bmod{28}$
$\left[30,30,3,1\right]$ $\left(1,19,10\right)$ $p\equiv 11,29\bmod{30}$
$\left[33,11,3,1\right]$ $\left(1,21,11\right)$ $p\equiv 2,8,17,29,32\bmod{33}$
$\left[40,40,2,1\right]$ $\left(1,19,20\right)$ $p\equiv 21,39\bmod{40}$
$\left[42,21,2,1\right]$ $\left(1,20,21\right)$ $p\equiv 5,17,41\bmod{42}$
Table 5. Cyclic genus $10$
5. Non-Cyclic Examples
Our program identified $17$ possible ramification types $z=[c_{0},c_{1},c_{\infty},s]$ for curves $X$ admitting an abelian cover $\phi:X\to\mathbf{P}^{1}$ branched at three points with $s>1$ such that the genus $g$ satisfies $6\leq g\leq 10$.
Remark 5.1.
For $5\leq g\leq 10$, we encountered three cases with no inertia type realizing the ramification type: ($g=6$) $z=[8,8,8,2]$; ($g=8$) $z=[12,12,4,2]$; and ($g=10$) $z=[12,12,12,2]$.
In this section, we analyze the 14 valid cases and prove a general result about a class of examples which occur for every genus.
Throughout, we use the following notation for quotients. Suppose $g\in G$ has prime order $\ell$. Let $X_{g}=X/\langle g\rangle$ be the quotient curve and $\phi_{g}:X\to X_{g}$ be the quotient cover. For $b\in\{0,1,\infty\}$, the inertia group $I_{b}$ is independent of the ramification point $\eta\in X$ above $b$ since $G$ is abelian. If $g\in I_{b}$, then $\phi_{g}$ is ramified with index $\ell$ at $\eta$. If $g\not\in I_{b}$, then $\phi_{g}$ is not ramified at $\eta$.
5.1. A useful lemma
Lemma 5.2.
Let $r\geq 2$. If $z=[2r,2r,2,2]$, then $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{0})\times\mathbf{Jac}(X_{1})$, where $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ is an abelian cover branched at $B=\{0,1,\infty\}$ for $i=1,2$. If $r$ is odd and $i=1,2$, or if $r$ is even and $i=1$, then $\phi_{i}$ has degree $2r$ with inertia type $a=(1,r,r-1)$. If $r$ is even, then $\phi_{0}$ has Galois group $\mathbf{Z}/r\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$ with inertia type $a=((1,0),(-1,-1),(0,1))$.
Proof.
By the Riemann–Hurwitz formula, $X$ has genus $r-1$. By Lemma 2.6, $G=\mathbf{Z}/2r\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. Without loss of generality, $a_{0}=(1,0)$, $a_{1}=(-1,1)$, and $a_{\infty}=(0,1)$. Since $a_{\infty}\in I_{\infty}$, the double cover $\phi_{a_{\infty}}:X\to X_{a_{\infty}}$ is branched at the $2r$ points above $\infty$. Thus $X_{a_{\infty}}$ has genus 00.
Let $K=\langle(r,0),(0,1)\rangle$. Note that $X/K$ also has genus 00, since it is a quotient of $X_{a_{\infty}}$. Write $X_{0}=X/\langle(r,0)\rangle$ and $X_{1}=X/\langle(r,1)\rangle$. By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{0})\times\mathbf{Jac}(X_{1})$.
For $g=(r,0)$ and $g=(r,1)$, note that $g\not\in I_{\infty}$; so $\phi_{g}$ is not ramified in the fiber over $\infty$. There are $2$ points of $X$ above each of 00 and $1$. Now $(r,0)\in I_{0}$ and $(r,1)\not\in I_{0}$. Also $(r,0)\in I_{1}$ if and only if $r$ is even, while $(r,1)\in I_{1}$ if and only if $r$ is odd.
Suppose $r$ is odd. By the Riemann–Hurwitz formula, $X_{0}$ and $X_{1}$ both have genus $(r-1)/2$. The quotient group $\overline{G}_{0}=G/\langle(r,0)\rangle$ is cyclic of order $2r$ with generator $\tau_{0}=(1,1)$. Then $a_{1}=-\tau_{0}$, $a_{\infty}\equiv r\tau_{0}\bmod(r,0)$, and $a_{0}=(r+1)\tau_{0}$. Thus the reduction of the inertia type in $\overline{G}_{0}$ is $(r+1,r,-1)\approx(1,r,r-1)$. The quotient group $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,0)=a_{0}$. We see that $a_{\infty}\equiv r\tau_{1}$ and $a_{1}\equiv(r-1)\tau_{1}$. Thus the $\mathbf{Z}/2r\mathbf{Z}$-cover $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ has inertia type $(1,r,r-1)$ for $i=1,2$.
Suppose $r$ is even. Then $\phi_{(r,1)}$ is unramified and $X_{1}$ has genus $r/2$; also $\phi_{(r,0)}$ is ramified at four points and $X_{0}$ has genus $(r/2)-1$. Also $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,1)=-a_{1}$. By a similar argument, $\phi_{1}$ has inertia type $(1,r,r-1)$. The inertia type for $\phi_{0}$ is found by working in $G/\langle(r,0)\rangle$. ∎
Remark 5.3.
Recall that $\mathbf{S}_{g}$ is the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. Using Lemma 5.2, for $r\geq 2$, one can show that the ramification type $z=[2r,2r,2,2]$ yields a curve $X$ of genus $r-1$ which is supersingular for all primes $p\equiv-1\bmod 2r$. This shows that the natural density of $\mathbf{S}_{g}$ is positive. We omit the details because this can also be proven using earlier work from [Yui80]. An unpublished study of supersingular diagonal curves (analyzed as quotients of the Fermat curve) also shows the density is positive [CMS].
5.2. Non-Cyclic Examples: Genus 6 through 10
The tables below contain the $14$ non-cyclic cases for $6\leq g\leq 10$. For most ramification types $z$, there is only one equivalence class of inertia type. In each case, we computed the decomposition of the Jacobian. The second column, titled decomposition, gives the data $(m_{i},\alpha_{i})$ indexed by $i\in I$. This indicates that there is an isogeny $\mathbf{Jac}(X)\sim\prod_{i\in I}\mathbf{Jac}(Z_{i})$, where there is a cyclic degree $m_{i}$ cover $\phi_{i}:Z_{i}\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ with inertia type $\alpha_{i}$. We use an addition sign to denote a product of Jacobians, an exponent to denote a product with multiplicity, and a subtraction sign to denote a quotient. From the decomposition, we used the Shimura–Taniyama method for cyclic covers implemented in our program to determine the congruence conditions on $p$, with $\mathrm{gcd}(p,|G|)=1$, for which $X$ is supersingular.
Lemma 5.4.
There are two ramification types for genus six curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
$z$ decomposition supersingular primes
$[5,5,5,5]$ $(5,(1,1,3))^{3}$ $p\equiv 2,3,4\bmod{5}$
$[14,14,2,2]$ $(14,(1,6,7))^{2}$ $p\equiv 3,5,13\bmod{14}$
Proof.
When $g=[5,5,5,5]$, then $a_{0}=(1,0)$, $a_{1}=(-1,-1)$, and $a_{\infty}=(0,1)$. By Theorem 2.12, $\mathbf{Jac}(X)$ is isogenous to the product of the Jacobians of six quotients of $X$. For $g=(1,0)$, $(0,1)$, and $(1,1)$, the cover $\phi_{g}$ is ramified at the five points above one point $b\in B$; thus $X_{g}$ has genus 00. For $g=(2,1)$, $(3,1)$, and $(4,1)$, the cover $\phi_{g}$ is not ramified and $X_{g}$ has genus two; the cover $\psi_{g}:X_{g}\to\mathbf{P}^{1}$ is a $\mathbf{Z}/5\mathbf{Z}$-cover branched at $B$; up to equivalence, there is only one inertia type for $\psi_{g}$, namely $(1,1,3)$.
The decomposition for $z=[14,14,2,2]$ follows from Lemma 5.2. ∎
Lemma 5.5.
There are three ramification types for genus seven curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
$z$ decomposition supersingular primes
$[9,9,3,3]$ $(3,(1,1,1))+(9,(1,3,5))+(9,(1,2,6))$ $p\equiv 2\bmod{3}$
$[12,6,4,2]$ $(12,(1,2,9))+(12,(1,3,8))$ $p\equiv 11\bmod{12}$
$[16,16,2,2]$ $(16,(1,7,8))+(8,(1,3,4))+(4,(1,1,2))$ $p\equiv 15\bmod{16}$
Proof.
Suppose $z=[9,9,3,3]$. Then $G=\mathbf{Z}/9\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ by Lemma 2.6. Then $a_{0}=(1,0)$, $a_{1}=(-1,-1)$, and $a_{\infty}=(0,1)$. Let $H\subset G$ be generated by $g_{0}=(3,0)$ and $g_{1}=(0,1)$. The cover $\phi_{(0,1)}$ is ramified at the nine points above $\infty$; thus $X_{(0,1)}$ has genus 00. By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{(3,0)})\times\mathbf{Jac}(X_{(3,1)})\times\mathbf{Jac}(X_{(3,2)})$. The cover $\phi_{(3,0)}$ is ramified at the three points above 00 and at the three points above $1$. Thus $X_{(3,0)}$ has genus $1$ and admits a cyclic degree $3$ cover of $\mathbf{P}^{1}$, with inertia type $(1,1,1)$.
The quotient group $G/\langle(3,2)\rangle$ is cyclic of order $9$, generated by $\tau=(1,1)=-a_{1}$. Then $a_{0}\equiv 7\tau$ and $a_{\infty}\equiv 3\tau$. So $\phi_{(3,2)}$ has inertia type $(7,3,-1)\approx(1,2,6)$. The quotient group $G/\langle(3,1)\rangle$ is cyclic of order $9$, generated by $\tau=(1,1)=-a_{1}$. Then $a_{0}\equiv 4\tau$ and $a_{\infty}\equiv 6\tau$. So $\phi_{(3,2)}$ has inertia type $(4,6,-1)\approx(1,3,5)$.
Suppose $z=[12,6,4,2]$. Then $G=\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. Then $a_{0}=(1,0)$, $a_{1}=(2,1)$, and $a_{\infty}=(9,1)$. Let $K=\langle(6,0),(0,1)\rangle$. The quotient $X_{(0,1)}$ has genus $4$ and is a $\mathbf{Z}/12\mathbf{Z}$ cover with inertia type $(1,2,9)$. The quotient $X_{(6,1)}$ has genus $3$ and is a $\mathbf{Z}/12\mathbf{Z}$ cover with inertia type $(1,3,8)$. The quotient $X_{(6,0)}$ has genus $2$ and is a $\mathbf{Z}/6\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$ cover with inertia type $((1,0),(2,1),(3,1))$. The quotient by $K$ is an elliptic curve $D$ which has an automorphism of order $3$. By Theorem 2.12, $\mathbf{Jac}(X_{(6,0)})\sim D^{2}$. Thus $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{(0,1)})\times\mathbf{Jac}(X_{(6,1)})$.
The case $[16,16,2,2]$ follows from Lemma 5.2 (used twice). ∎
Lemma 5.6.
There are two ramification types for genus eight curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
$z$ decomposition supersingular primes
$[10,10,10,2]$ $(10,(1,1,8))+(10,(1,3,6))^{2}-(5,(1,1,3))$ $p\equiv 3,7,9\bmod{10}$
$[10,10,10,2]$ $(10,(1,1,8))+(10,(1,2,7))^{2}-(5,(1,1,3))$ $p\equiv 3,7,9\bmod{10}$
$[18,18,2,2]$ $(18,(1,8,9))^{2}$ $p\equiv 5,11,17\bmod{18}$
Proof.
When $z=[10,10,10,2]$, there are two possibilities for $a$, up to equivalence: $a=((1,0),(1,1),(8,1))$ or $a=((1,0),(2,1),(7,1))$. We apply Theorem 2.12 using the Klein-four group $K$. The quotients by the three involutions are $\mathbf{Z}/10\mathbf{Z}$-covers of $\mathbf{P}^{1}$ branched at $B$. The quotient $X/K$ has genus $2$ and admits a $\mathbf{Z}/5\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with $a=(1,1,3)$.
The case $[18,18,2,2]$ follows from Lemma 5.2. ∎
We determine enough information to show that abelian non-cyclic covers produce no new primes for supersingular curves of genus $9$.
Lemma 5.7.
There are three ramification types for genus nine curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
$z$ supersingular primes
$[8,8,4,4]$ necessary condition $p\equiv 7\bmod 8$
$[12,12,6,2]$ necessary condition $p\equiv 11\bmod 12$
$[20,20,2,2]$ $p\equiv 19\bmod{20}$
Proof.
For $z=[8,8,4,4]$, $G=\mathbf{Z}/8\mathbf{Z}\times\mathbf{Z}/4\mathbf{Z}$ and $a=((1,0),((7,3),(0,1))$. We consider the quotient by $g=(4,2)$. Since $(4,2)$ is not in $I_{0}$, $I_{1}$, or $I_{\infty}$, the cover $\phi_{(4,2)}$ is unramified. Thus $X_{(4,2)}$ has genus $5$ and Galois group $\overline{G}=(\mathbf{Z}/8\mathbf{Z}\times\mathbf{Z}/4\mathbf{Z})/\langle(4,2)\rangle$. Note that $\overline{G}\simeq\mathbf{Z}/8\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$, with generators $\tau=(1,1)$ and $\sigma=(2,1)$. The inertia type reduces to $7\tau+\sigma$, $7\tau$, $2\tau+\sigma$. After multiplying by $-\tau$ and rearranging, this is Case (6) from $g=5$. By Proposition 3.3, $X_{(4,2)}$ is supersingular if and only if $p\equiv 7\bmod 8$.
For $z=[12,12,6,2]$, the group is $\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$. Then $a=((1,0),(1,1),(10,1))$. The quotient by $(0,1)$ is unramified, so $X_{(0,1)}$ has genus $5$. It admits a cyclic degree 12 cover of $\mathbf{P}^{1}$ branched at $B$ with inertia type $(1,1,10)$. This is Case (7) from $g=5$. By Proposition 3.6, $X$ is supersingular only when $p\equiv 11\bmod 12$.
The case $[20,20,2,2]$ decomposes as $(20,(1,9,10))+(10,(1,4,5))^{2}$ by Lemma 5.2. ∎
Lemma 5.8.
There are four ramification types for genus ten curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$. They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
$z$ decomposition supersingular primes
$[6,6,6,6]$ partial information below $p\equiv 5\bmod 6$
$[9,9,9,3]$ $(9,(1,1,7))^{4}-(3,(1,1,1))^{2}$ $p\equiv 2\bmod{3}$
$[12,12,3,3]$ $(12,(1,3,8)^{2})+(12,(1,4,7))$ $p\equiv 11\bmod{12}$
$[22,22,2,2]$ $(20,(1,9,10))^{2}$ $p\equiv 11,19\bmod{20}$
Proof.
If $z=[6,6,6,6]$, then $G=\mathbf{Z}/6\mathbf{Z}\times\mathbf{Z}/6\mathbf{Z}$ by Lemma 2.6 and we may take $a=((1,0),(0,1),(5,5))$. Take $H=\langle(2,0),(0,2)\rangle\simeq(\mathbf{Z}/3\mathbf{Z})^{2}$. For $g=(2,0)$, $(0,2)$, or $(2,2)$, the cover $\phi_{g}$ is ramified at 6 points, so $X_{g}$ has genus $2$. Also $X_{g}$ has an automorphism of order $6$.
For $g=(2,4)$, the cover $\phi_{g}$ is unramified and so $X_{(2,4)}$ has genus $4$. The group $\overline{G}=G/\langle(2,4)\rangle$ is isomorphic to $\mathbf{Z}/6\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$, generated by $\tau=(1,1)$ and $\sigma=(1,2)$. Note that $(a_{0},a_{1},a_{\infty})$ reduces to $(2\tau+\sigma,5\tau+\sigma,5\tau)\approx((2,1),(5,1),(5,0))$. In the Klein four action on $X_{(2,4)}$, each of the three involutions fixes two points, so the quotient by an involution gives a curve of genus $2$ that has an automorphism of order $6$.
By [Igu60, Section 8, page 645], a curve of genus $2$ with an automorphism of order $6$ is isomorphic to $y^{2}=x(x-1)(x+1)(x-2)(x-(1/2))$. By [IKO86, Proposition 1.11], this curve is supersingular if and only if $p\equiv 5\bmod 6$. Applying Theorem 2.12 twice (to the quotient of $X_{2,4}$ by the Klein four action and to the quotient of $X$ by $H$) shows that $X$ is supersingular if and only if $p\equiv 5\bmod 6$.
If $z=[9,9,9,3]$, then $G=\mathbf{Z}/9\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ and $a=((1,0),(1,1),(7,2))$. Let $H=(\mathbf{Z}/3\mathbf{Z})^{2}$. The quotient $X_{(3,0)}$ and the quotient $X/H$ are both elliptic curves, having an automorphism of order $3$. Each of $X_{(3,1)}$, $X_{(3,2)}$ and $X_{(0,1)}$ has genus $4$, and is a $\mathbf{Z}/9\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with inertia type $\bar{a}\approx(1,1,7)$. Then use Theorem 2.12.
If $z=[12,12,3,3]$, then $G=\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ and $a=((1,0),(11,2),(0,1))$. Let $H=(\mathbf{Z}/3\mathbf{Z})^{2}$. The quotients $X_{(0,1)}$ and $X_{H}$ both have genus 00. By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{(4,0)})\times\mathbf{Jac}(X_{(4,1)})\times\mathbf{Jac}(X_{(4,2)})$. Then $X_{(4,0)}$ is a $\mathbf{Z}/4\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with $a=((1,0),(3,2),(0,1))$, which is equivalent to a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(1,3,8)$. For $X_{(4,1)}$, we note that $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,1)\rangle$ is generated by $\tau=(1,2)$. So $X_{(4,1)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,3\tau,4\tau)\approx(1,3,8)$. Similarly, $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,2)\rangle$ is generated by $\tau=(1,1)$ and $X_{(4,2)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,11\tau,8\tau)\approx(1,4,7)$.
The case $[22,22,2,2]$ follows from Lemma 5.2. ∎
6. Conclusions
Theorem 6.1.
For $5\leq g\leq 10$: the primes $p$ for which there exists a supersingular curve $X$ of genus $g$ having an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ are:
  1. (1)
    for $g=5$, any $p$ such that $p\equiv QNR\bmod 11$ and/or
    $p\equiv 7,11,19,23,31,41,47,59,71,79,83,89,91,101,103,107,119\bmod 120.$
    Thus $\delta_{5}=25/32$.
  2. (2)
    for $g=6$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv 2,3,4\bmod 5$; $p\equiv 3,5,6\bmod 7$; $p\not\equiv 1,3,9\bmod 13$; $p\equiv 15\bmod 16$; and $p\equiv 13\bmod 24$. Thus $\delta_{6}=507/512$.
  3. (3)
    for $g=7$, any $p$ such that $p\equiv 2\bmod 3$ and/or $p\equiv 3\bmod 4$. Thus $\delta_{7}=3/4$.
  4. (4)
    for $g=8$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv 2,3,4\bmod 5$; $p\not\equiv 1\bmod 17$; and $p\not\equiv 1,15\bmod 32$. Thus $\delta_{8}=1023/1024$.
  5. (5)
    for $g=9$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv 3\bmod 4$; $p\equiv 3,5,6\bmod 7$; and $p\equiv QNR\bmod 19$. Thus $\delta_{9}=15/16$.
  6. (6)
    for $g=10$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv QNR\bmod 11$; and $p\not\equiv 1\bmod 40$. Thus $\delta_{10}=63/64$.
Proof.
For $g=5$, this follows from Propositions 3.1, 3.3, and 3.6. For $6\leq g\leq 10$, this follows from the tables in Sections 4 and 5. ∎

### v2

@@
 Abstract.
-For a curve which admits an abelian cover of the projective line branched at three points, we study when its reduction to positive characteristic is supersingular.
-Using the method of Shimura and Taniyama, we give a complete classification when the genus of the curve is at most 10.
-The natural density of the set of primes for which this construction yields a supersingular curve is larger than expected.
+For a curve $X$ which is an abelian cover of the projective line branched at three points, we study when its reduction to positive characteristic is supersingular.
+Using the method of Shimura and Taniyama, we give a complete classification of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular when the genus of $X$ is at most $10$.
+The natural density of this set of primes is larger than expected.
 2020 MSC primary: 11G20, 11M38, 14G10, 14H10, 14H40. Secondary 11G10, 14G15, 14H37, 14H30.
+Key words and phrases:
+curve, Jacobian, abelian variety, positive characteristic, supersingular, Frobenius, Newton polygon, automorphism
 Pries was supported by NSF grant DMS-22-00418.
-We thank the organizers of the RIMS conference on Theory and Applications of Supersingular Curves and Supersingular Abelian Varieties, II.
-We thank Everett Howe, Wanlin Li, and Elena Mantovan for helpful conversations.
+This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University We thank the organizers of the RIMS conference on Theory and Applications of Supersingular Curves and Supersingular Abelian Varieties, II.
+We thank Everett Howe, Wanlin Li, Elena Mantovan, and Darren Schmidt for helpful conversations, and the referee for helpful comments.
 1. Introduction
-Suppose $k$ is an algebraically closed field of prime characteristic $p$.
+Let $k$ be an algebraically closed field of prime characteristic $p$.
 An elliptic curve over $k$ is supersingular if it has no points of order $p$.
-A smooth projective curve $X$ of genus $g$ over $k$ is supersingular if its Jacobian is isogenous to a product of $g$ supersingular elliptic curves.
-One good method for constructing supersingular curves is to study a curve $X$ that is an abelian cover of the projective line $\mathbf{P}^{1}$ branched at exactly three points.
-The degree and ramification indices of the cover determine the genus $g$ of $X$.
-Such curves are quotients of a Fermat curve and their Jacobians have complex multiplication.
+A smooth projective curve $X$ of genus $g$ over $k$ is supersingular if its Jacobian ${\mathbf{Jac}}(X)$ is isogenous to a product of $g$ supersingular elliptic curves.
+One good method for constructing supersingular curves is to study families of curves that are abelian covers of the projective line $\mathbf{P}^{1}$.
+The degree and ramification indices of the cover determine the genus $g$.
+When the curve $X$ is branched at exactly three points, it is a quotient of a Fermat curve and ${\mathbf{Jac}}(X)$ has complex multiplication.
 This makes it easier to study the action of Frobenius on the cohomology [Yui80] and hence determine whether the reduction of $X$ modulo $p$ is supersingular.
-Recently, we were curious which pairs $(g,p)$ occur for supersingular curves in this context.
-Our motivation was to show that a new construction of supersingular curves of genus $5$ [BP] applied for infinitely many new primes.
+For example, Re studied supersingular quotients of the Fermat curve with the aim of producing supersingular curves of many genera in fixed characteristic [Re08, Re09].
+Recently, we were curious about an analogous question: given $g$, for which primes $p$ do these constructions yield a supersingular curve of genus $g$ in characteristic $p$?
+Our motivation was to show that a new construction of supersingular curves [BP] yields supersingular curves of genus $5$ in characteristic $p$ for infinitely many new primes $p$.
 We wrote this paper because we could not find a reference that states what can be determined using this method.
-Let $\mathbf{S}_{g}$ be the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at exactly three points.
-In this paper, we determine $\mathbf{S}_{g}$ for $5\leq g\leq 10$.
-The main result is Theorem 6.1, which builds on data found in Sections 3, 4 and 5.
-We summarize it as follows.
+Let $\mathbf{S}_{g}$ be the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k=\overline{\mathbf{F}}_{p}$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at exactly three points.
+For $5\leq g\leq 10$, we determine the primes for which there exists a supersingular curve of this type in Sections 3, 4 and 5.
+This yields the main result of this paper, Theorem 6.1, which states exactly which primes are in $\mathbf{S}_{g}$, for $5\leq g\leq 10$, in terms of congruence conditions.
+From these congruence conditions, we determine the natural density of the set $\mathbf{S}_{g}$, which we state as the following corollary.
 Corollary 1.1.
-For $5\leq g\leq 10$, here is the natural density $\delta_{g}$ of the set $\mathbf{S}_{g}$:
+For $5\leq g\leq 10$, the natural density $\delta_{g}$ of the set $\mathbf{S}_{g}$ is:
 […]
-The cases $g=9,10$ are especially interesting.
-The codimension of the supersingular locus in the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties of dimension $g$ is $g(g+1)/2-\lfloor g^{2}/4\rfloor$.
-For $g\geq 9$, this exceeds the dimension of the moduli space $\mathcal{M}_{g}$ of smooth curves of genus $g$.
-For this reason, the existence of a smooth supersingular curve of genus $g$ is geometrically unexpected for $g\geq 9$.
-Work of Yui [Yui80] implies that $\delta_{g}$ is positive for all $g$, which can also be deduced from our analysis, see Remark 5.3.
-For the proof, we use Ekedahl’s work on [Eke87, page 173] to classify the degrees and ramification indices of abelian covers $X\to\mathbf{P}^{1}$ branched at 3 points with fixed genus $g$.
-When $g=4,5$, Ekedahl determines conditions on $p$ for which $X$ is superspecial (meaning its Jacobian is isomorphic to a product of supersingular elliptic curves).
-We work instead with the Shimura–Taniyama formula [Tat71, Section 5], a method used also in [LMPT19b], to determine the set of primes $p$ for which the reduction is supersingular.
-We use SageMath to enumerate all the possibilities for genus $5\leq g\leq 10$.
-We restrict to the case $g\geq 5$ and $p$ odd because of earlier work: for $g=1$ by Deuring [Deu41]; for $g=2$ by Serre [Ser83, Théorème 3], also [IKO86, Proposition 3.1]; for $g=3$ by Oort [Oor91, Theorem 5.12], also [Ibu93, Theorem 1]; and for $g=4$ by Harashita, Kudo, and Senda [KHS20, Corollary 1.2], also [Pri24, Theorem 1.1].
-For $p=2$, for all $g$, see the work of van der Geer and van der Vlugt [vdGvdV95].
+[…]
+Work of Yui [Yui80] implies that $\delta_{g}$ is positive for all $g\geq 1$, which can also be deduced from our analysis, see Remark 5.3.
+For the proof, we use Ekedahl’s work [Eke87, p.
+173] to classify the degrees and ramification indices of abelian covers $X\to\mathbf{P}^{1}$ branched at $3$ points, such that $X$ has genus $g$.
+When $g=4,5$, Ekedahl determines conditions on $p$ for which $X$ is superspecial (meaning ${\mathbf{Jac}}(X)$ is isomorphic to a product of supersingular elliptic curves).
+We determine the set of primes $p$ for which the reduction is supersingular using the Shimura–Taniyama formula [Tat71, Section 5], a method also used in [LMPT19b].
+As explained in Section 2.4, this method exploits the structure of complex multiplication on ${\mathbf{Jac}}(X)$ to determine its Newton polygon.
+We use SageMath to enumerate all abelian covers of $\mathbf{P}^{1}$ ramified at three points with genus $g$ satisfying $5\leq g\leq 10$.
+We restrict to the case $g\geq 5$ and $p$ odd because of previous constructions of supersingular curves: for $g=1$ by Deuring [Deu41]; for $g=2$ by Serre [Ser83, Théorème 3], also [IKO86, Proposition 3.1]; for $g=3$ by Oort [Oor91, Theorem 5.12], also [Ibu93, Theorem 1]; for $g=4$ by Harashita, Kudo, and Senda [KHS20, Corollary 1.2], also [Pri, Theorem 1.1]; and for $p=2$ and arbitrary $g$, by van der Geer and van der Vlugt [vdGvdV95].
+Remark 1.2.
+Many mathematicians studied supersingular curves of genus $g\geq 5$ in characteristic $p$ for various pairs $(g,p)$.
+Here is an incomplete list: Hermitian curves [Han92], Artin-Schreier curves [vdGvdV92, vdGvdV95, BHM+16], cyclic covers of $\mathbf{P}^{1}$ [Re08, Re09, LMPT19b, LMPT19a]; and trigonal curves [Kud19, KH22].
 2. Abelian covers branched at three points
@@
 We suppose that the Galois group $G$ of $\phi$ is abelian and that $\phi$ is tamely ramified.
-2.1. The degree and ramification indices
+2.1. Ramification and Inertia Types
 For $b\in B$, let $I_{b}$ be the inertia group above $b$ and let $c_{b}=|I_{b}|$.
@@
 Let $r=\mathrm{gcd}(c_{0},c_{1},c_{\infty})$. For distinct $i,j\in B$, let $e_{i,j}=\mathrm{gcd}(c_{i},c_{j})/r$.
-Any two of $e_{0,1}$, $e_{0,\infty}$, and $e_{1,\infty}$ are relatively prime and $e_{0,1}\geq e_{0,\infty}\geq e_{1,\infty}$.
+Ekedahl [Eke87, page 173] verified that $e_{0,1}$, $e_{0,\infty}$, and $e_{1,\infty}$ are pairwise relatively prime, that $e_{0,1}\geq e_{0,\infty}\geq e_{1,\infty}$; and that
+(1) $c_{0}=re_{0,1}e_{0,\infty},\quad c_{1}=re_{0,1}e_{1,\infty},\quad c_{\infty}=re_{0,\infty}e_{1,\infty}.$
 Lemma 2.3.
 [Eke87, page 173] With notation as above:
-(2.1) $2g-2=s(re_{0,1}e_{0,\infty}-(1+(e_{0,1}/e_{1,\infty})+(e_{0,\infty}/e_{1,\infty})).$
+(2) $2g-2=s(re_{0,1}e_{0,\infty}-(1+(e_{0,1}/e_{1,\infty})+(e_{0,\infty}/e_{1,\infty})).$
 Proof.
@@
 $=$ $s(c_{0}-(1+(c_{0}/c_{1})+(c_{0}/c_{\infty}))).$
-The result follows since $c_{0}=re_{0,1}e_{0,\infty}$, $c_{1}=re_{0,1}e_{1,\infty}$, and $c_{\infty}=re_{0,\infty}e_{1,\infty}$. ∎
-We identify $G$ with $\mathrm{Gal}(\phi)$ and define a refinement of the ramification type.
+The result follows from (1). ∎
+A cover $\phi$ can be described via the numerical data of the ramification type.
+However, there are also group-theoretic constraints on the ramification.
+To state these, we identify $G$ with $\mathrm{Gal}(\phi)$ and define the inertia type.
 Definition 2.4.
-See e.g., [Bou01, Definitions 3.2 & 4.1].
-The inertia type $a$ is the triple $a=(a_{0},a_{1},a_{\infty})$, where $a_{b}$ denotes the canonical generator for $I_{b}$, found from the action of the inertia group on a local parameter at a ramification point above $b$.
+Following Bouw [Bou01, Definitions 3.2 & 4.1], the inertia type of $\phi$ is the triple $a=(a_{0},a_{1},a_{\infty})$, where $a_{b}$ denotes the canonical generator for $I_{b}$, found from the action of inertia on a local parameter at a ramification point above $b$.
 There are three constraints on the inertia type: $|a_{b}|=c_{b}$ for $b\in B$; $a_{0},a_{1},a_{\infty}$ generate $G$; and $a_{0}a_{1}a_{\infty}=\mathrm{Id}_{G}$.
 This implies that $s$ divides $c_{1}$ and $s$ divides $c_{\infty}$.
-Not every ramification type $z=[c_{0},c_{1},c_{\infty},s]$ has an inertia type that satisfies these constraints.
+Given a tuple $z=[c_{0},c_{1},c_{\infty},s]$ satisfying the numerical constraints for a ramification type, it is apriori possible that there is no tuple $a$ satisfying the group-theoretic constraints for the inertia type.
 See Remark 5.1 for some examples.
-Given $z$, we can change the labeling of branch points whose inertia groups have the same order; this allows us to permute these entries of the inertia type.
-Also, we can change the identification of $G$ with $\mathrm{Gal}(\phi)$; this modifies each entry of the inertia type by the action of some $\sigma\in\mathrm{Aut}(G)$.
-Two inertia types are equivalent, denoted $\approx$, if they are in the same orbit under these actions.
+Given an inertia type, we can change the labeling of branch points whose inertia groups have the same order, permuting those entries of the inertia type.
+Also, we can change the identification of $G$ with $\mathrm{Gal}(\phi)$; this modifies each entry of the inertia type by some $\sigma\in\mathrm{Aut}(G)$.
+Two inertia types $a$ and $a^{\prime}$ are equivalent, denoted $a\approx a^{\prime}$, if they are in the same orbit under these actions.
 2.2. Cyclic covers branched at three points
-In this section, we suppose $\phi$ is cyclic of order $m$, and identify $G=\mathbf{Z}/m\mathbf{Z}$.
+Suppose $\phi:X\to{\mathbf{P}}^{1}$ is cyclic of order $m$, and identify $G=\mathbf{Z}/m\mathbf{Z}$.
 The inertia type $(a_{0},a_{1},a_{\infty})$ is defined up to multiplication by $(\mathbf{Z}/m\mathbf{Z})^{*}$.
@@
 An affine equation for the cover is
-$X:y^{m}=x^{a_{0}}(1-x)^{a_{1}},$
+(3) $X:y^{m}=x^{a_{0}}(1-x)^{a_{1}},$
 with $G$ acting by multiplication by $m$th roots of unity on $y$.
@@
   • if $1\leq j<m$, then $f_{j}\in\{0,1\}$, and $f(j)+f(m-j)=1$ provided $d_{j}$, the order of $j$ modulo $m$, does not divide $a_{b}$ for any $b\in B$.
+2.3. Newton polygons
+If $X$ is a smooth curve over $k$, we describe the Newton polygon of its Jacobian ${\mathbf{Jac}}(X)$ in terms of the decomposition of the $p$-divisible group ${\mathbf{Jac}}(X)[p^{\infty}]$.
+Definition 2.6.
+For non-negative integers $c$ and $d$ with ${\mathrm{gcd}}(c,d)=1$, fix a $p$-divisible group $G_{c,d}$ of codimension $c$, dimension $d$, and thus height $c+d$.
+The slope of $G_{c,d}$ is $\lambda=d/(c+d)$, with multiplicity $c+d$.
+Let $G_{\lambda}=G_{c,d}$.
+Suppose $A$ is a principally polarized abelian variety of dimension $g$ over $k$.
+By the Dieudonné–Manin classification [Man63], the $p$-divisible group $A[p^{\infty}]$ is isogenous to $\oplus_{\lambda}G_{\lambda}^{m_{\lambda}}$ for some multi-set of non-negative integers $m_{\lambda}$.
+This multi-set defines a symmetric Newton polygon of height $2g$ with integer break-points.
+As a short-hand, we write $\nu(A)=\oplus_{\lambda}G_{\lambda}^{m_{\lambda}}$ and call both $\nu(A)$ and its visual representation the Newton polygon of $A$.
+The Newton polygon of a curve $X$ over $k$ is the Newton polygon of ${\mathbf{Jac}}(X)$.
+A curve $X$ of genus $g$ is ordinary if $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{g}$ and is supersingular if $\nu({\mathbf{Jac}}(X))=G_{1,1}^{g}$.
+2.4. Newton polygons for cyclic covers branched at three points
+Suppose $\phi:X\to\mathbf{P}^{1}$ is a cyclic degree $m$ cover which is branched at three points.
+Assuming that $p$ does not divide $m$, we describe how to compute the Newton polygon of ${\mathbf{Jac}}(X)$ from the signature type of $\phi$, using a formula due to Shimura and Taniyama, see [LMPT19b, §2,3] for details.
+This implicitly relies on $X$ being the reduction of a curve over $\mathbf{Q}$ whose Jacobian has complex multiplication.
+Notation 2.7.
+Assume that $p$ does not divide $m$.
+Let $R$ denote the set of orbits of the multiplication by $p$ map [×$[\times p]$ on the set $\{j\in\mathbf{Z}/m\mathbf{Z}:j\neq 0\}$.
+For each orbit $r\in R$, the additive order $d_{j}$ of $j$ modulo $m$ is the same for all values $j$ in $r$ and we denote it by $d_{r}$.
+For $\epsilon\in\{0,1\}$, let
+[…]
+For $r\in R$ such that $d_{r}$ does not divide $a_{b}$ for $b\in B$, set $\alpha_{r}=\#(r\cap S_{1})$.
+In the context of Notation 2.7, the set $R$ is in one-to-one correspondence with the set of primes of the group ring $\mathbf{Q}[\mu_{m}]$ above $p$.
+This is because $\mathbf{Q}[\mu_{m}]$ is unramified at $p$; the elements of $\mathbf{Z}/m\mathbf{Z}$ can be identified with homomorphisms from $\mathbf{Q}[\mu_{m}]$ to $\mathbf{C}$; and there is a natural action of Frobenius on the set of homomorphisms, see [LMPT19b, §2.1] for details.
+The $p$-divisible group ${\mathbf{Jac}}(X)[p^{\infty}]$ canonically decomposes into pieces ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ indexed by the orbits $r\in R$.
+Theorem 2.8 (Shimura–Taniyama formula).
+[Tat71, Section 5] The only slope of the Newton polygon of ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ is $\lambda_{r}:=\alpha_{r}/\#r$.
+In particular, if every orbit is self-dual, meaning invariant under negation modulo $m$, then $X$ is supersingular.
+Example 2.9.
+Let $m=8$ and $p\equiv 3\bmod 4$. The orbits are $r_{1}=\{1,3\}$, $r_{2}=\{2,6\}$, $r_{5}=\{5,7\}$, and $r_{4}=\{4\}$.
+If $a=(1,1,6)$, then $\phi$ has signature $f=(1,1,1,0,0,0,0)$.
+Thus $S_{1}=\{1,2,3\}$.
+Then $\lambda_{r_{1}}=1$, $\lambda_{r_{2}}=1/2$, and $\lambda_{r_{5}}=0$.
+Thus $\nu(\mathbf{Jac}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$.
+If $a=(1,5,2)$, then $\phi$ has signature $f=(1,1,0,0,1,0,0)$. Thus $S_{1}=\{1,2,5\}$ and $X$ is supersingular.
+Remark 2.10.
+There is another method to determine whether $X$ is supersingular by Yui [Yui80] and Aoki [Aok08] using $p$-adic valuations of Jacobi sums.
+By [Aok08, Theorem 1.1], $X$ is supersingular if and only if either (i) $p^{i}\equiv-1\bmod m$ for some $i$; or (ii) there is an equivalence (as defined in Section 2.1) of inertia types $a\approx(1,-p^{i},p^{i}-1)$ for some $i$ such that $d=\mathrm{gcd}(p^{i}-1,m)>1$ and $p^{j}\equiv-1\bmod m/d$ for some $j$.
 2.5. Non-cyclic Galois groups and inertia types
-The following determines the Galois group in all the cases with $s>1$ for $5\leq g\leq 10$.
+Suppose $\phi$ is abelian but not cyclic, of order $m$. The following lemma determines the Galois group of $\phi$ in all the cases with $s>1$ for $5\leq g\leq 10$.
 Lemma 2.11.
@@
 By hypothesis, there is a prime $\ell$ dividing all three of $s=c_{0}s/c_{0}$, $c_{0}s/c_{1}$, and $c_{0}s/c_{\infty}$.
-If $G$ was cyclic, $\ell$ would divide each of $m,a_{0},a_{1},a_{\infty}$, contradicting the fact that $a_{0},a_{1},a_{\infty}$ generate $G$.
+If $G$ was cyclic, $\ell$ would divide each of $m=c_{0}s,a_{0},a_{1},a_{\infty}$, contradicting the fact that $a_{0},a_{1},a_{\infty}$ generate $G$.
 Since $G$ is generated by two elements $a_{0},a_{1}$, this implies that $G\simeq\mathbf{Z}/r_{1}\mathbf{Z}\times\mathbf{Z}/r_{2}\mathbf{Z}$ where $r_{1}r_{2}=c_{0}s$ and $r_{2}\mid r_{1}$.
@@
 For $s>1$, we denote elements of $G=\mathbf{Z}/c_{0}\mathbf{Z}\times\mathbf{Z}/s\mathbf{Z}$ as an ordered pair.
-So each $a_{b}$ in the inertia type $a=(a_{0},a_{1},a_{\infty})$ is written as $a_{b}=(a_{b,1},a_{b,2})$.
-To choose a representative for the equivalence class of $a$, we take the natural inclusion of $G$ in $V=(\mathbf{Z}/c_{0}\mathbf{Z})^{2}$, and use that $\mathrm{Aut}(V)$ acts transitively on pairs of linearly independent elements of order $c_{0}$.
-Without loss of generality, we choose $a_{0}=(1,0)$.
+Each $a_{b}$ in the inertia type $a=(a_{0},a_{1},a_{\infty})$ is written as $a_{b}=(a_{b,1},a_{b,2})$.
+To choose a representative for the equivalence class of $a$, we take the natural inclusion of $G$ in $V=(\mathbf{Z}/c_{0}\mathbf{Z})^{2}$.
+Since $\mathrm{Aut}(V)$ acts transitively on pairs of linearly independent elements of order $c_{0}$, we may choose $a_{0}=(1,0)$ without loss of generality.
 When $c_{\infty}=s$, we also choose $a_{\infty}=(0,1)$.
-2.4. Newton polygons for cyclic covers branched at three points
-Suppose $p$ is a prime not dividing $m$, the degree of the cyclic cover $\phi$.
-We consider the reductions of $X$ and ${\mathbf{Jac}}(X)$ modulo $p$.
-The Newton polygon of ${\mathbf{Jac}}(X)$ is an invariant that measures the $p$-adic valuations of the eigenvalues of Frobenius.
-In this section, we explain how to compute the Newton polygon from the signature type of $\phi$.
-Definition 2.7.
-For non-negative integers $c$ and $d$ with ${\mathrm{gcd}}(c,d)=1$, fix a $p$-divisible group $G_{c,d}$ of codimension $c$, dimension $d$, and thus height $c+d$.
-The slope of $G_{c,d}$ is $\lambda=d/(c+d)$, with multiplicity $c+d$.
-Let $G_{\lambda}=G_{c,d}$.
-Suppose $A$ is a principally polarized abelian variety of dimension $g$ over $k$, By the Dieudonné–Manin classification, the $p$-divisible group $A[p^{\infty}]$ is isogenous to $\oplus_{\lambda}G_{\lambda}^{m_{\lambda}}$ for some multi-set of non-negative integers $m_{\lambda}$.
-This multi-set defines a symmetric Newton polygon $\nu(A)$ of height $2g$ with integer break-points.
-Remark 2.8.
-Suppose $A$ is defined over $\mathbf{F}_{0}$, a finite field of characteristic $p$.
-Let $W(\mathbf{F}_{0})$ denote the Witt vector ring of $\mathbf{F}_{0}$.
-Consider the action of Frobenius $\varphi$ on the crystalline cohomology group $H^{1}_{\rm cris}(A/W(\mathbf{F}_{0}))$.
-There exists an integer $n$ such that $\varphi^{n}$, the composite of $n$ Frobenius actions, is a linear map.
-The Newton polygon $\nu(A)$ is the multi-set of rational numbers $\lambda$ such that $n\lambda$ are the valuations at $p$ of the eigenvalues of Frobenius for this action.
-We sometimes describe the Newton polygon of a curve $X$ in terms of the decomposition of ${\mathbf{Jac}}(X)[p^{\infty}]$.
-For example, a curve of genus $g$ is ordinary if its $p$-divisible group is $(G_{0,1}\oplus G_{1,0})^{g}$ and is supersingular if its $p$-divisible group is $G_{1,1}^{g}$.
-Notation 2.9.
-Let $R$ denote the set of orbits of the multiplication by $p$ map [×$[\times p]$ on the set $\{j\in\mathbf{Z}/m\mathbf{Z}:j\neq 0\}$.
-For each orbit $r\in R$, the additive order $d_{j}$ of $j$ modulo $m$ is the same for all values $j$ in $r$ and we denote it by $d_{r}$.
-For $\epsilon\in\{0,1\}$, let
-[…]
-For $r\in R$ such that $d_{r}$ does not divide $a_{b}$ for $b\in B$, set $\alpha_{r}=\#(r\cap S_{1})$.
-The set $R$ is in one-to-one correspondence with the set of primes of the group ring $\mathbf{Q}[\mu_{m}]$ above $p$.
-The $p$-divisible group ${\mathbf{Jac}}(X)[p^{\infty}]$ canonically decomposes into pieces ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ indexed by the orbits $r\in R$.
-By the Shimura–Taniyama formula [Tat71, Section 5], the only slope of the Newton polygon of ${\mathbf{Jac}}(X)[p^{\infty}]_{r}$ is $\lambda_{r}:=\alpha_{r}/\#r$.
-In particular, if every orbit is self-dual, meaning invariant under negation modulo $m$, then $X$ is supersingular.
-Example 2.10.
-Let $m=8$ and $p\equiv 3\bmod 4$. The orbits are $r_{1}=\{1,3\}$, $r_{2}=\{2,6\}$, $r_{5}=\{5,7\}$, and $r_{4}=\{4\}$.
-If $a=(1,1,6)$, then $\phi$ has signature $f=(1,1,1,0,0,0,0)$.
-Thus $S_{1}=\{1,2,3\}$.
-Then $\lambda_{r_{1}}=1$, $\lambda_{r_{2}}=1/2$, and $\lambda{r_{5}}=0$.
-Thus $\nu(\mathbf{Jac}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$.
-If $a=(1,5,2)$, then $\phi$ has signature $f=(1,1,0,0,1,0,0)$. Thus $S_{1}=\{1,2,5\}$ and $X$ is supersingular.
-Remark 2.11.
-One can also determine whether $X$ is supersingular using a method of Yui [Yui80] and Aoki [Aok08] that relies on $p$-adic valuations of Jacobi sums.
-By [Aok08, Theorem 1.1], $X$ is supersingular if and only if either (i) $p^{i}\equiv-1\bmod m$ for some $i$; or (ii) $a\approx(1,-p^{i},p^{i}-1)$ for some $i$ such that $d=\mathrm{gcd}(p^{i}-1,m)>1$ and $p^{j}\equiv-1\bmod m/d$ for some $j$.
 2.6. Decomposition from the Kani–Rosen theorem
@@
 Then there is an isogeny:
-(2.3) $\mathrm{Jac}(X)^{t-1}\times\mathrm{Jac}(X/G)^{n_{G}}\sim\prod_{i=1}^{t}\mathrm{Jac}(X/H_{i})^{n_{i}}.$
+(5) $\mathrm{Jac}(X)^{t-1}\times\mathrm{Jac}(X/G)^{n_{G}}\sim\prod_{i=1}^{t}\mathrm{Jac}(X/H_{i})^{n_{i}}.$
+Example 2.13.
 Suppose $G\simeq(\mathbf{Z}/2\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3}\subset G$ be the three subgroups of order two.
-Then (2.3) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{2}\sim\prod_{i=1}^{3}\mathrm{Jac}(X/H_{i})$.
+Then (5) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{2}\sim\prod_{i=1}^{3}\mathrm{Jac}(X/H_{i})$.
 Suppose $G\simeq(\mathbf{Z}/3\mathbf{Z})^{2}$ and let $H_{1},H_{2},H_{3},H_{4}\subset G$ be the four subgroups of order three.
-Then (2.3) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{3}\sim\prod_{i=1}^{4}\mathrm{Jac}(X/H_{i})$.
+Then (5) simplifies to $\mathrm{Jac}(X)\times\mathrm{Jac}(X/G)^{3}\sim\prod_{i=1}^{4}\mathrm{Jac}(X/H_{i})$.
 2.7. Enumerating and analyzing abelian covers
-Using SageMath, we automated the search for ramification data producing abelian covers $\phi:X\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ whose genus lies in a specified range.
-We outline the approach, which depends on whether $s=1$ and $s>1$.
-When $s=1$, then $\phi$ is cyclic and the program determines possible inertia types.
-Without loss of generality, we identified the generator of $I_{0}$ with $1\in G$ and so $a_{0}=1$.
-Then the program analyzes the congruence classes of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular, using the Shimura–Taniyama formula as in Section 2.4.
-For each case when $s>1$, Lemma 2.6 implies that $G$ is non-cyclic and that $G\simeq(\mathbf{Z}/c_{0}\mathbf{Z})\times(\mathbf{Z}/s\mathbf{Z})$.
-For every $z$, we determined the possible inertia types $a$.
-Without loss of generality, we identified the generator of $I_{0}$ with $a_{0}=(1,0)$.
-In the small number of non-cyclic cases with $5\leq g\leq 10$, we determined by hand a decomposition of $\mathbf{Jac}(X)$ into Jacobians of curves that are cyclic covers of $\mathbf{P}^{1}$ branched at $B$, using the Kani–Rosen method as in Section 2.5.
-This allowed us to find the supersingular primes.
+Using SageMath, we automated the search for abelian covers $\phi:X\to\mathbf{P}^{1}$ over $\mathbf{Q}$ branched at $B=\{0,1,\infty\}$ whose genus lies in a specified range.
+We then partially automated the determination of the set of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular.
+We outline the approach here: the code is on github [BP25].
+We begin by listing all possible $e_{0,1},e_{0,\infty},e_{1,\infty},r,s$ which satisfy equation (2) for $g$ in the specified range. These determine the potential ramification types of covers.
+For a potential ramification type with $s=1$, the cover $\phi$ is cyclic and $G=\mathbf{Z}/m\mathbf{Z}$.
+We look for inertia types compatible with the ramification type.
+Without loss of generality, $a_{0}=1$, and we search for $a_{1}$ and $a_{\infty}$ with $|a_{1}|=c_{1}$, $|a_{\infty}|=c_{\infty}$, and $a_{0}+a_{1}+a_{\infty}=0\mod{m}$.
+Then (3) gives a cover with that inertia type.
+Finally, the program determines the congruence classes of primes $p$ for which the reduction of $X$ modulo $p$ is supersingular, using the Shimura–Taniyama formula found in Theorem 2.8.
+When $s>1$, Lemma 2.11 implies that $\phi$ is non-cyclic and that $G\simeq(\mathbf{Z}/c_{0}\mathbf{Z})\times(\mathbf{Z}/s\mathbf{Z})$.
+Again, we look for inertia types compatible with the ramification type.
+Without loss of generality, we identify the potential generator of $I_{0}$ with $a_{0}=(1,0)$ and search for $a_{1}$ and $a_{\infty}$ satisfying the constraints listed after Definition 2.4.
+Note that an inertia type realizing a potential ramification type may not exist, in which case there is no cover $\phi$ realizing that ramification type; see Remark 5.1 for examples.
+In the small number of non-cyclic cases with $5\leq g\leq 10$, for a curve $X$ with inertia type $a$, we determine by hand a decomposition of $\mathbf{Jac}(X)$ into Jacobians of curves that are cyclic covers of $\mathbf{P}^{1}$ branched at $B$, using the Kani–Rosen method as in Section 2.6.
+The program then gives us the primes for which the reduction of $X$ is supersingular.
 In Section 3, we give a complete analysis of the genus $5$ situation to illustrate the methods.
-In particular, abelian covers of $\mathbf{P}^{1}$ branched at three points do not produce supersingular curves of genus $5$ for every prime $p\equiv 3\bmod 4$.
-Section 4 contains the results of our program [BP25] for cyclic cases when $6\leq g\leq 10$.
+Section 4 contains the results of our program for cyclic cases when $6\leq g\leq 10$.
 Section 5 treats the non-cyclic cases when $6\leq g\leq 10$.
 3. Refining Ekedahl’s Results for genus five
-Suppose $\phi:X\to\mathbf{P}^{1}$ is a tamely ramified abelian cover branched at $B=\{0,1,\infty\}$.
+Suppose $\phi:X\to\mathbf{P}^{1}$ is a tamely ramified abelian cover over $k$ branched at $B=\{0,1,\infty\}$.
 In [Eke87, page 173], Ekedahl enumerates ways $\phi$ can ramify so that $X$ has genus $5$.
@@
 The superspecial condition is equivalent to the Frobenius operator being trivial on $H^{1}(X,\mathcal{O}_{X})$.
-For each of the seven cases, Ekedahl determines the primes $p$ for which $X$ is superspecial, thus proving that there exists a superspecial smooth curve of genus $5$ which is an abelian cover of $\mathbf{P}^{1}$ branched at $3$ points if and only if $p\equiv-1\bmod 11,12,15,20,8$ or $p\equiv 11\bmod 15$.
+For each of the seven cases, Ekedahl determines the primes $p$ for which $X$ is superspecial, thus proving that there exists a superspecial smooth curve of genus $5$ which is an abelian cover of $\mathbf{P}^{1}$ branched at $3$ points if and only if $p\equiv-1\bmod 8,11,12,15,20$ or $p\equiv 11\bmod 15$.
 If $X$ is superspecial, then it is supersingular, but the converse is not always true.
@@
 Proposition 3.1.
-If $\phi:X\to\mathbf{P}^{1}$ is a cyclic cover branched at $3$ points and $X$ has genus $5$, then the ramification type $z$ for $\phi$ is one of the following:
-  1. (1)
-    $z=[11,11,11,1]$. Then $X$ is supersingular if and only if $p$ is a QNR modulo $11$.
-  2. (2)
-    $z=[15,15,3,1]$. Then $X$ is supersingular if and only if $p\equiv-1,11\bmod 15$.
-  3. (3)
-    $z=[22,11,2,1]$. Then $X$ is supersingular if and only if $p$ is a QNR modulo $11$.
-  4. (4)
-    $z=[12,12,6,1]$. Then $X$ is supersingular if and only if $p\equiv-1\bmod 12$.
-  5. (5)
-    $z=[20,20,2,1]$. Then $X$ is supersingular if and only if $p\equiv-1,11\bmod 20$.
+If $\phi:X\to\mathbf{P}^{1}$ is a cyclic cover branched at $3$ points and $X$ has genus $5$, then its ramification type $z$ and set of supersingular primes are one of:
+Case z supersingular primes
+$1$ $[11,11,11,1]$ $p\equiv\text{QNR}\bmod 11$
+$2$ $[15,15,3,1]$ $p\equiv-1,11\bmod 15$
+$3$ $[22,11,2,1]$ $p\equiv\text{QNR}\bmod 11$
+$4$ $[12,12,6,1]$ $p\equiv-1\bmod 12$
+$5$ $[20,20,2,1]$ $p\equiv-1,11\bmod 20$
 Proof.
@@
 Recall that $R$ denotes the set of orbits of ×$\times p$ on $\mathbf{Z}/m\mathbf{Z}$.
-In each situation, if $p^{i}\equiv-1\bmod m$ for some $i$, then each orbit is self-dual and so $X$ is supersingular.
+If $p^{i}\equiv-1\bmod m$ for some $i$, then each orbit is self-dual and so $X$ is supersingular.
 When $m$ is composite, there can be additional cases where $\alpha_{r}/\#r=1/2$ for each $r\in R$.
-We give some additional details in each case.
+Here are some details.
   1. (1)
-    Note that $a\approx(1,k,-(1+k))$ for some $1\leq k\leq 9$. The result is proven as one case of [LMPT19b, Theorem 5.1].
+    This is one case of [LMPT19b, Theorem 5.1].
   2. (2)
@@
 By the Shimura–Taniyama formula, $X$ is supersingular (but not superspecial).
-∎
 3.2. Non-cyclic cases for genus five
-Suppose $\phi:X\to\mathbf{P}^{1}$ is an abelian non-cyclic cover branched at $B=\{0,1,\infty\}$ and $X$ has genus $5$.
-Ekedahl shows there are two options for the ramification type: either (6) $z=[8,8,4,2]$ or (7) $z=[12,12,2,2]$.
-In both cases, we determine the congruence conditions on $p$ for which $X$ is supersingular.
+Suppose $\phi:X\to\mathbf{P}^{1}$ is an abelian non-cyclic cover over $k$ branched at $B=\{0,1,\infty\}$.
+When $X$ has genus $5$, Ekedahl shows there are two options for the ramification type: either (6) $z=[8,8,4,2]$ or (7) $z=[12,12,2,2]$.
 3.2.1. Case (6)
@@
 Lemma 3.2.
-When $z=[8,8,4,2]$, then:
-  1. (1)
-    the quotient of $X$ by $H$ is isomorphic to $E_{1728}$;
+When $z=[8,8,4,2]$, then $X/H$ is isomorphic to $E_{1728}$. Also:
   1. (1)
@@
 The $G$-cover $X\to{\mathbf{P}}^{1}$ has: $2$ points above 00 with $I_{0}=\langle(1,0)\rangle$; $2$ points above $1$ with $I_{1}=\langle(1,1)\rangle$; and $4$ points above $\infty$ with $I_{\infty}=\langle(6,1)\rangle$.
-  1. (1)
 Note that $G/H\simeq\mathbf{Z}/4\mathbf{Z}$.
 So $\psi:X/H\to{\mathbf{P}}^{1}$ is a $\mathbf{Z}/4\mathbf{Z}$-cover branched at $B$.
-The intersection of $H$ with each of $I_{0}$, $I_{1}$, and $I_{\infty}$ has order $2$.
-By the Riemann–Hurwitz formula, $X/H$ has genus $1$.
-Since $\psi$ is totally ramified above 00 and $\infty$, this implies that $X/H$ is isomorphic to $E_{1728}$.
+The intersection of $H$ with each of $I_{0}$, $I_{1}$, and $I_{\infty}$ has order $2$, so $X/H$ has genus $1$.
+Since $\psi$ is totally ramified above 00 and $1$, this implies that $X/H\cong E_{1728}$.
   1. (1)
@@
 Since $(4,0)$ is in the inertia group at every point of $X$ above 00, $1$, and $\infty$, the double cover $X\to X/{H_{1}}$ is branched at $8$ points.
-By the Riemann–Hurwitz formula, $X/{H_{1}}$ is an elliptic curve, which we denote $E^{\prime}$.
+So $X/{H_{1}}$ is an elliptic curve, which we denote $E^{\prime}$.
 The double cover $E^{\prime}\to E_{1728}$ is unramified and is thus a $2$-isogeny.
 Let $W$ be the quotient of $X$ by $H_{2}:=\langle(0,1)\rangle$.
-Since $H_{2}\cap I_{b}$ is trivial for each $b\in B$, the double cover $X\to W$ is unramified.
-By the Riemann–Hurwitz formula, $W$ has genus $3$.
+Since $H_{2}\cap I_{b}$ is trivial for each $b\in B$, the double cover $X\to W$ is unramified and $W$ has genus $3$.
 The same holds for the quotient $W^{\prime}$ of $X$ by $H_{3}:=\langle(4,1)\rangle$.
@@
 The result follows since the category of abelian varieties up to isogeny is semi-simple, and ${\mathbf{Jac}}(E_{1728})\sim{\mathbf{Jac}}(E^{\prime})$.
-∎
 Proposition 3.3.
-Suppose $z=[8,8,4,2]$. Then $X$ is supersingular if and only if $p\equiv 7\bmod 8$. If $p\equiv 3\bmod 8$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
+Suppose $z=[8,8,4,2]$. Then $X$ is supersingular if and only if $p\equiv 7\bmod 8$. If $p\equiv 3\bmod 8$, then $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
 Proof.
-By Lemma 3.2(i), $E_{1728}$ is a quotient of $X$.
+By Lemma 3.2, $E_{1728}$ is a quotient of $X$.
 We restrict to the case $p\equiv 3\bmod 4$ which is equivalent to $E_{1728}\sim E^{\prime}$ being supersingular.
-In this case, $X$ is supersingular if and only if $W$ and $W^{\prime}$ are both supersingular, by Lemma 3.2(iii).
+In this case, $X$ is supersingular if and only if $W$ and $W^{\prime}$ are both supersingular, by Lemma 3.2.
 The cover $W\to\mathbf{P}^{1}$ has Galois group $\mathbf{Z}/8\mathbf{Z}$ and is branched above $B$ with inertia type $\overline{a}=(1,1,6)$.
@@
 We compute its inertia type $\overline{a}^{\prime}=(5z,z,2z)\approx(1,5,2)$.
-If $p\equiv 7\bmod 8$, then every orbit of ×$\times p$ on $\{1\leq j<8\}$ is self-dual, and so both $W$ and $W^{\prime}$ are supersingular by the Shimura–Taniyama formula.
+If $p\equiv 7\bmod 8$, then every orbit of ×$\times p$ on $\{1\leq j<8\}$ is self-dual, and so both $W$ and $W^{\prime}$ are supersingular by Theorem 2.8.
 Suppose $p\equiv 3\bmod 8$.
-By Example 2.10, the curve $W^{\prime}$ is supersingular since $\overline{a}^{\prime}\approx(1,5,2)$, while the curve $W$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$ since $\overline{a}=(1,1,6)$.
-Thus ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})$ has two slopes of 00 and $1$, and $X$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
+By Example 2.9, the curve $W^{\prime}$ is supersingular since $\overline{a}^{\prime}\approx(1,5,2)$, while $\nu(\mathbf{Jac}(W))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}$ since $\overline{a}=(1,1,6)$.
+Thus ${\mathbf{Jac}}(X)\times{\mathbf{Jac}}(E_{1728})$ has two slopes of 00 and $1$, and $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
 ∎
@@
     The curve $Z$ has an action by $G/\langle(6,1)\rangle\simeq\mathbf{Z}/12\mathbf{Z}$, which is generated by $w=(1,1)$.
-The inertia type $a=((1,0),(11,1),(0,1))$ reduces to $\overline{a}=(7w,11w,6w)\approx(5,1,6)$.
-There is an order $3$ automorphism $\sigma$ on $Z$ that fixes two points.
-The quotient of $Z$ by $\sigma$ is an elliptic curve with an automorphism of order $4$ that fixes two points.
+The inertia type $a$ reduces to $\overline{a}=(7w,11w,6w)\approx(5,1,6)$.
+The curve $Z$ has an order $3$ automorphism $\sigma$ that fixes two points and $Z/\langle\sigma\rangle$ is an elliptic curve with an automorphism of order $4$ that fixes two points.
 Thus $Z$ is a $\mathbf{Z}/3\mathbf{Z}$-cover of $E_{1728}$.
@@
     The curve $W$ admits an action of $\mathbf{Z}/6\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$.
-Each of the automorphisms $(3,0)$ and $(3,1)$ fixes exactly two points of $W$, so $W$ admits two double covers of an elliptic curve.
-In both cases, this elliptic curve is $E_{0}$ because it has an automorphism of order $3$ which fixes three points.
+The automorphisms $(3,0)$ and $(3,1)$ each fix two points of $W$, so $W$ admits two double covers of an elliptic curve, which must be $E_{0}$ because it has an automorphism of order $3$ which fixes three points.
 Since $(3,0)$ and $(3,1)$ generate a Klein-four group, ${\mathbf{Jac}}(W)\sim E_{0}^{2}$ by Theorem 2.12.
-∎
 Proposition 3.6.
@@
 Then $X$ is supersingular if and only if $p\equiv 11\bmod 12$.
-If $p\equiv 5\bmod 12$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{3}\oplus G_{1,1}^{2}$.
-If $p\equiv 7\bmod 12$, then ${\mathbf{Jac}}(X)$ has Newton polygon $(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
+If $p\equiv 5\bmod 12$, then $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{3}\oplus G_{1,1}^{2}$.
+If $p\equiv 7\bmod 12$, then $\nu({\mathbf{Jac}}(X))=(G_{0,1}\oplus G_{1,0})^{2}\oplus G_{1,1}^{3}$.
 Proof.
@@
 By Lemma 3.5, ${\mathbf{Jac}}(X)\sim{\mathbf{Jac}}(Z)\times E_{0}^{2}$.
-The curve $Z$ is supersingular by the Shimura–Taniyama formula since the orbits of ×$\times p$ on $\mathbf{Z}/12\mathbf{Z}$ are self-dual.
+The curve $Z$ is supersingular by Theorem 2.8, since the orbits of ×$\times p$ on $\mathbf{Z}/12\mathbf{Z}$ are self-dual.
 This completes the proof of the first claim.
@@
 When $p\equiv 5\bmod 12$, then $E_{0}$ is supersingular.
-Also the orbits of ×$\times p$ are $\{1,5\}$, $\{7,11\}$, $\{3\}$ and $\{9\}$.
-By the Shimura–Taniyama formula, the Newton polygon of $Z$ is $(G_{0,1}\oplus G_{1,0})^{3}$.
+Also, the orbits of ×$\times p$ are $\{1,5\}$, $\{7,11\}$, $\{3\}$ and $\{9\}$.
+By Theorem 2.8, $\nu(\mathbf{Jac}(Z))=(G_{0,1}\oplus G_{1,0})^{3}$.
 When $p\equiv 7\bmod 12$, then $E_{0}$ is ordinary.
-Also the orbits of ×$\times p$ are $\{1,7\}$, $\{5,11\}$, and $\{3,9\}$.
-By the Shimura–Taniyama formula, $Z$ is supersingular.
+Also, the orbits of ×$\times p$ are $\{1,7\}$, $\{5,11\}$, and $\{3,9\}$.
+By Theorem 2.8, $Z$ is supersingular.
 ∎
 4. Cyclic Examples Genus 6 through 10
-Our program computed the following data for curves $X$, of genus $6\leq g\leq 10$, that admit a degree $m$ cyclic cover $\phi:X\to\mathbf{P}^{1}$ branched at three points: the ramification type $z$; the possible inertia types $a$, up to equivalence; and the odd primes $p$ with $\mathrm{gcd}(p,m)=1$ such that $X$ is supersingular.
-The results are listed in Tables 1-5
+Our program computed this data for curves $X$ over $k$, of genus $6\leq g\leq 10$, that occur as a degree $m$ cyclic cover $\phi:X\to\mathbf{P}^{1}$ branched at three points: the ramification types $z$; the inertia types $a$, up to equivalence; and the odd primes $p$ with $\mathrm{gcd}(p,m)=1$ for which the reduction of $X$ modulo $p$ is supersingular.
+The results are in Tables 1-5.
 $z$ $a$ supersingular primes
@@
 5. Non-Cyclic Examples
-Our program identified $17$ possible ramification types $z=[c_{0},c_{1},c_{\infty},s]$ for curves $X$ admitting an abelian cover $\phi:X\to\mathbf{P}^{1}$ branched at three points with $s>1$ such that the genus $g$ satisfies $6\leq g\leq 10$.
+Our program identified $14$ ramification types $z=[c_{0},c_{1},c_{\infty},s]$, with $s>1$, for which there is a valid inertia type, for an abelian cover $\phi:X\to\mathbf{P}^{1}$ branched at three points such that the genus $g$ of $X$ satisfies $6\leq g\leq 10$.
+In this section, we analyze the supersingular primes $p$ for these covers, assuming throughout that $p\nmid c_{0}s$.
+We also prove a general result about a class of examples which occur for every genus.
 Remark 5.1.
 For $5\leq g\leq 10$, we encountered three cases with no inertia type realizing the ramification type: ($g=6$) $z=[8,8,8,2]$; ($g=8$) $z=[12,12,4,2]$; and ($g=10$) $z=[12,12,12,2]$.
-In this section, we analyze the 14 valid cases and prove a general result about a class of examples which occur for every genus.
-Throughout, we use the following notation for quotients.
-Suppose $g\in G$ has prime order $\ell$.
-Let $X_{g}=X/\langle g\rangle$ be the quotient curve and $\phi_{g}:X\to X_{g}$ be the quotient cover.
+Here is some notation.
+If $g\in G$ has prime order $\ell$, let $X_{g}=X/\langle g\rangle$ be the quotient curve and $\phi_{g}:X\to X_{g}$ be the quotient cover.
 For $b\in\{0,1,\infty\}$, the inertia group $I_{b}$ is independent of the ramification point $\eta\in X$ above $b$ since $G$ is abelian.
@@
 Let $r\geq 2$.
-If $z=[2r,2r,2,2]$, then $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{0})\times\mathbf{Jac}(X_{1})$, where $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ is an abelian cover branched at $B=\{0,1,\infty\}$ for $i=1,2$.
-If $r$ is odd and $i=1,2$, or if $r$ is even and $i=1$, then $\phi_{i}$ has degree $2r$ with inertia type $a=(1,r,r-1)$.
-If $r$ is even, then $\phi_{0}$ has Galois group $\mathbf{Z}/r\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$ with inertia type $a=((1,0),(-1,-1),(0,1))$.
+If $z=[2r,2r,2,2]$, then $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{1})\times\mathbf{Jac}(X_{2})$, where $\phi_{i}:X_{i}\to\mathbf{P}^{1}$ is an abelian cover branched at $B=\{0,1,\infty\}$ for $i=1,2$.
+If $r$ is odd and $i=1,2$, or if $r$ is even and $i=1$, then $\phi_{i}$ has group $\mathbf{Z}/2r\mathbf{Z}$ with $a=(1,r,r-1)$.
+If $r$ is even, then $\phi_{2}$ has group $\mathbf{Z}/r\mathbf{Z}\times\mathbf{Z}/2\mathbf{Z}$ with $a=((1,0),(-1,-1),(0,1))$.
 Proof.
@@
 Let $K=\langle(r,0),(0,1)\rangle$.
-Note that $X/K$ also has genus 00, since it is a quotient of $X_{a_{\infty}}$.
-Write $X_{0}=X/\langle(r,0)\rangle$ and $X_{1}=X/\langle(r,1)\rangle$.
-By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{0})\times\mathbf{Jac}(X_{1})$.
+Since $X/K$ is a quotient of $X_{a_{\infty}}$, it also has genus 00.
+Write $X_{1}=X/\langle(r,1)\rangle$ and $X_{2}=X/\langle(r,0)\rangle$.
+By Theorem 2.12, $\mathbf{Jac}(X)\sim\mathbf{Jac}(X_{1})\times\mathbf{Jac}(X_{2})$.
 For $g=(r,0)$ and $g=(r,1)$, note that $g\not\in I_{\infty}$; so $\phi_{g}$ is not ramified in the fiber over $\infty$.
@@
 Suppose $r$ is odd.
-By the Riemann–Hurwitz formula, $X_{0}$ and $X_{1}$ both have genus $(r-1)/2$.
-The quotient group $\overline{G}_{0}=G/\langle(r,0)\rangle$ is cyclic of order $2r$ with generator $\tau_{0}=(1,1)$.
+Then $X_{1}$ and $X_{2}$ both have genus $(r-1)/2$.
+The quotient group $\overline{G}_{2}=G/\langle(r,0)\rangle$ is cyclic of order $2r$ with generator $\tau_{0}=(1,1)$.
 Then $a_{1}=-\tau_{0}$, $a_{\infty}\equiv r\tau_{0}\bmod(r,0)$, and $a_{0}=(r+1)\tau_{0}$.
-Thus the reduction of the inertia type in $\overline{G}_{0}$ is $(r+1,r,-1)\approx(1,r,r-1)$.
+Thus the reduction of the inertia type in $\overline{G}_{2}$ is $(r+1,r,-1)\approx(1,r,r-1)$.
 The quotient group $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,0)=a_{0}$.
@@
 Suppose $r$ is even.
-Then $\phi_{(r,1)}$ is unramified and $X_{1}$ has genus $r/2$; also $\phi_{(r,0)}$ is ramified at four points and $X_{0}$ has genus $(r/2)-1$.
+Then $\phi_{(r,1)}$ is unramified and $X_{1}$ has genus $r/2$; also $\phi_{(r,0)}$ is ramified at four points and $X_{2}$ has genus $(r/2)-1$.
 Also $\overline{G}_{1}=G/\langle(r,1)\rangle$ is cyclic of order $2r$ with generator $\tau_{1}=(1,1)=-a_{1}$.
 By a similar argument, $\phi_{1}$ has inertia type $(1,r,r-1)$.
-The inertia type for $\phi_{0}$ is found by working in $G/\langle(r,0)\rangle$.
+The inertia type for $\phi_{2}$ is found by working in $G/\langle(r,0)\rangle$.
 ∎
 Remark 5.3.
-Recall that $\mathbf{S}_{g}$ is the set of primes $p$ for which there exists a supersingular curve $X$ of genus $g$ defined over $k$ such that $X$ is an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$.
-Using Lemma 5.2, for $r\geq 2$, one can show that the ramification type $z=[2r,2r,2,2]$ yields a curve $X$ of genus $r-1$ which is supersingular for all primes $p\equiv-1\bmod 2r$.
+By Lemma 5.2, for $r\geq 2$, the ramification type $z=[2r,2r,2,2]$ yields a curve $X$ of genus $r-1$ which is supersingular for all primes $p\equiv-1\bmod 2r$.
 This shows that the natural density of $\mathbf{S}_{g}$ is positive.
-We omit the details because this can also be proven using earlier work from [Yui80].
-An unpublished study of supersingular diagonal curves (analyzed as quotients of the Fermat curve) also shows the density is positive [CMS].
+We omit the details because this can also be proven using [Yui80] or (unpublished) work in [CMS].
 5.2. Non-Cyclic Examples: Genus 6 through 10
-The tables below contain the $14$ non-cyclic cases for $6\leq g\leq 10$.
-For most ramification types $z$, there is only one equivalence class of inertia type.
-In each case, we computed the decomposition of the Jacobian.
-The second column, titled decomposition, gives the data $(m_{i},\alpha_{i})$ indexed by $i\in I$.
-This indicates that there is an isogeny $\mathbf{Jac}(X)\sim\prod_{i\in I}\mathbf{Jac}(Z_{i})$, where there is a cyclic degree $m_{i}$ cover $\phi_{i}:Z_{i}\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ with inertia type $\alpha_{i}$.
+The tables below contain the following information for the $14$ non-cyclic cases for $6\leq g\leq 10$: the ramification type $z$; the decomposition of the Jacobian; the congruence conditions on $p$, assuming $\mathrm{gcd}(p,|G|)=1$, for which the curve $X$ is supersingular.
+The decomposition of the Jacobian is given by the data $(m_{i},\alpha_{i})$ indexed by $i\in I$, meaning that there is an isogeny $\mathbf{Jac}(X)\sim\prod_{i\in I}\mathbf{Jac}(Z_{i})$, where there is a cyclic degree $m_{i}$ cover $\phi_{i}:Z_{i}\to\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ with inertia type $\alpha_{i}$.
 We use an addition sign to denote a product of Jacobians, an exponent to denote a product with multiplicity, and a subtraction sign to denote a quotient.
-From the decomposition, we used the Shimura–Taniyama method for cyclic covers implemented in our program to determine the congruence conditions on $p$, with $\mathrm{gcd}(p,|G|)=1$, for which $X$ is supersingular.
 Lemma 5.4.
-There are two ramification types for genus six curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$.
-They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
+Here are the supersingular primes for genus six curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
 $z$ decomposition supersingular primes
@@
 Lemma 5.5.
-There are three ramification types for genus seven curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$.
-They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
+Here are the supersingular primes for genus seven curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
 $z$ decomposition supersingular primes
@@
 Lemma 5.6.
-There are two ramification types for genus eight curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$.
-They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
+Here are the supersingular primes for genus eight curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
 $z$ decomposition supersingular primes
@@
 Lemma 5.7.
-There are three ramification types for genus nine curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$.
-They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
+Here are the supersingular primes for genus nine curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
 $z$ supersingular primes
@@
 By Proposition 3.6, $X$ is supersingular only when $p\equiv 11\bmod 12$.
-The case $[20,20,2,2]$ decomposes as $(20,(1,9,10))+(10,(1,4,5))^{2}$ by Lemma 5.2. ∎
+By Lemma 5.2, $[20,20,2,2]$ decomposes as $(20,(1,9,10))+(10,(1,4,5))^{2}$. ∎
 Lemma 5.8.
-There are four ramification types for genus ten curves that admit a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$.
-They yield supersingular curves under the following conditions on $p\nmid c_{0}s$.
+Here are the supersingular primes for genus ten curves realized as a non-cyclic abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$:
 $z$ decomposition supersingular primes
@@
 $[12,12,3,3]$ $(12,(1,3,8)^{2})+(12,(1,4,7))$ $p\equiv 11\bmod{12}$
-$[22,22,2,2]$ $(20,(1,9,10))^{2}$ $p\equiv 11,19\bmod{20}$
+$[22,22,2,2]$ $(22,(1,10,11))^{2}$ $p\equiv 7,13,17,19,21\bmod{22}$
 Proof.
@@
 By [IKO86, Proposition 1.11], this curve is supersingular if and only if $p\equiv 5\bmod 6$.
-Applying Theorem 2.12 twice (to the quotient of $X_{2,4}$ by the Klein four action and to the quotient of $X$ by $H$) shows that $X$ is supersingular if and only if $p\equiv 5\bmod 6$.
+Applying Theorem 2.12 twice shows that $X$ is supersingular if and only if $p\equiv 5\bmod 6$.
 If $z=[9,9,9,3]$, then $G=\mathbf{Z}/9\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$ and $a=((1,0),(1,1),(7,2))$.
 Let $H=(\mathbf{Z}/3\mathbf{Z})^{2}$.
-The quotient $X_{(3,0)}$ and the quotient $X/H$ are both elliptic curves, having an automorphism of order $3$.
+The quotients $X_{(3,0)}$ and $X/H$ are both elliptic curves, having an automorphism of order $3$.
 Each of $X_{(3,1)}$, $X_{(3,2)}$ and $X_{(0,1)}$ has genus $4$, and is a $\mathbf{Z}/9\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with inertia type $\bar{a}\approx(1,1,7)$.
@@
 Then $X_{(4,0)}$ is a $\mathbf{Z}/4\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z}$-cover of $\mathbf{P}^{1}$ branched at $B$ with $a=((1,0),(3,2),(0,1))$, which is equivalent to a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(1,3,8)$.
-For $X_{(4,1)}$, we note that $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,1)\rangle$ is generated by $\tau=(1,2)$.
-So $X_{(4,1)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,3\tau,4\tau)\approx(1,3,8)$.
+Because $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,1)\rangle$ is generated by $\tau=(1,2)$, $X_{(4,1)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,3\tau,4\tau)\approx(1,3,8)$.
 Similarly, $(\mathbf{Z}/12\mathbf{Z}\times\mathbf{Z}/3\mathbf{Z})/\langle(4,2)\rangle$ is generated by $\tau=(1,1)$ and $X_{(4,2)}$ is a $\mathbf{Z}/12\mathbf{Z}$-cover with inertia type $(5\tau,11\tau,8\tau)\approx(1,4,7)$.
@@
 Theorem 6.1.
-For $5\leq g\leq 10$: the primes $p$ for which there exists a supersingular curve $X$ of genus $g$ having an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ are:
+The primes $p$ for which there exists a supersingular curve $X$ of genus $5\leq g\leq 10$ having an abelian cover of $\mathbf{P}^{1}$ branched at $B=\{0,1,\infty\}$ are:
   1. (1)
-    for $g=5$, any $p$ such that $p\equiv QNR\bmod 11$ and/or
-    $p\equiv 7,11,19,23,31,41,47,59,71,79,83,89,91,101,103,107,119\bmod 120.$
-    Thus $\delta_{5}=25/32$.
+    for $g=5$, any $p\equiv 7,11,19,23,31,41,47,59,71,79,83,89,91,101,103,107,119\bmod 120$ and/or $p\equiv QNR\bmod 11$. Thus $\delta_{5}=25/32$.
   2. (2)
@@
   6. (6)
-    for $g=10$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv QNR\bmod 11$; and $p\not\equiv 1\bmod 40$. Thus $\delta_{10}=63/64$.
+    for $g=10$, any $p$ satisfying at least one of the conditions: $p\equiv 2\bmod 3$; $p\equiv QNR\bmod 11$; $p\not\equiv 1,11,31\bmod 40$; and $p\equiv 3,19,27\pmod{28}$. Thus $\delta_{10}=31/32$.
 Proof.